### v1

# AL-Bench: A Benchmark for Automatic Logging

Abstract.
Logging, the practice of inserting log statements into source code, is critical for improving software reliability. Recently, language model-based techniques have been developed to automate log generation based on input code. Although these methods demonstrate promising results in isolated evaluations, their effectiveness diminishes when applied to ad-hoc low-quality data and code similarity-based evaluation methods. We consider a comprehensive evaluation benchmark should include (1) a high-quality, diverse, and large-scale dataset, (2) an assessment of the compilability of the code with inserted log statements, and (3) a runtime log-oriented evaluation method.
To this end, this paper introduces AL-Bench, a comprehensive benchmark designed specifically for automatic logging tools. AL-Bench includes a high-quality, diverse dataset collected from 10 widely recognized projects with varying logging requirements and introduces a novel dynamic evaluation approach. Different from the evaluation in existing logging papers, AL-Bench assesses both the compilability of the code with inserted log statements and the quality of the logs generated by them during runtime, which we believe can better reflect the effectiveness of logging techniques in practice. AL-Bench reveals significant limitations in the state-of-the-art tools The codes with log statements generated by the state-of-the-art tools fail to compile in 20.1%-83.6% cases. In addition, even the best-performing tool did not achieve high similarity between the runtime logs produced by the generated log statements and the ground-truth log statements, demonstrating a 0.213 cosine similarity. The results reveal substantial opportunities to further enhance the development of automatic logging tools.
Logging, Software Maintenance, Logging Benchmark
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Software and its engineering
1. introduction
As software grows in size and complexity, logging has become increasingly essential to ensuring software reliability (He et al., 2021; Chen and Jiang, 2021). Logging means writing log statements into the source code, which generate runtime logs that record valuable information for a range of downstream tasks such as anomaly detection (Zhang et al., 2019; Lou et al., 2010; He et al., 2016; Du et al., 2017; Zhang et al., 2019), fault diagnosis (Zou et al., 2016), root cause analysis (Amar and Rigby, 2019; He et al., 2018b; Lin et al., 2016), and program verification (Ding et al., 2015; Shang et al., 2013). The effectiveness of these downstream tasks heavily relies on the quality of the software logs (He et al., 2018a). Therefore, appropriate logging is essential to capture critical behaviors during software operation (Yuan et al., [n. d.]).
Figure 1. An example of logging statement generation. Logging statement generation can be separated as three parts: determining the position, selecting the level, and specifying the message.
Numerous logging tools have been developed to assist software developers by automatically suggesting log statements based on provided code snippets (Mastropaolo et al., 2022; Liu et al., 2021; Xu et al., 2024; Xie et al., 2023). As illustrated in Figure 1, automatic logging typically includes three steps: (1) determining the position, (2) deciding the verbosity level, and (3) specifying the message to be recorded. Leveraging the advanced text generation capabilities of large language models (LLMs) (Floridi and Chiriatti, 2020; Raffel et al., 2020), LANCE is the first end-to-end tool to integrate positioning, level selection, and message generation. UniLog (Xu et al., 2024) employed a warm-up and in-context learning (ICL) strategy to enhance performance. FastLog (Xie et al., 2023) improved the generation efficiency while maintaining precision. LEONID (Mastropaolo et al., 2023), based on LANCE, combined with deep learning and information retrieval technologies to enhance performance. SCLogger (Li et al., 2024) adapted static analysis to extend the context for the code snippet. These studies typically evaluate performance using ad-hoc data splitting from the entire dataset, focusing on metrics such as the accuracy of log statement components, including position, level, and message. Machine translation metrics, such as BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004), are also employed to evaluate the quality of generated log messages. Although these evaluations offer valuable insights into logging tool performance, the use of low-quality data and incomplete assessments undermines the reliability of the results.
First, loose standards collection and inappropriate clean strategy compromise the quality of evaluation data and effectiveness in assessing the performance. Previous evaluation datasets were typically created by splitting the entire dataset. The data selection rules of the entire dataset are commonly loose standards to ensure sufficient data for training. These criteria fail to ensure the quality and consistency of the data. Moreover, to accommodate the limitations of the tools, they filtered out all instances exceeding 512 tokens, ignoring the long code snippets that are commonly encountered in real-world development environments. This approach undermines the overall effectiveness and real-world applicability of the evaluation results, as it does not accurately reflect the true complexity of software projects.
Second, the current evaluation method does not verify whether the generated log statements are compilable. Generating the compilable log statements is the first requirement when applying automatic logging tools in practice. To relieve developers from the heavy effort required to design and maintain log statements (Chen and Jiang, 2019, 2017), the basic requirement is to ensure that our tool can be seamlessly integrated into the DevOps process without introducing additional errors that require extra debugging effort from developers. Current evaluation methods merely focus on whether each component of log statements (i.e., position, verbosity level, message) matches the ground truth but cannot assess whether the generated log statements might introduce compilation errors such as wrong code format or using undefined variables. Evaluating the compilability of predicted log statements reflects the effectiveness and reliability of logging tools, which are essential for practical use.
Third, the evaluation method cannot evaluate the quality of runtime logs generated by the predicted log statements. Current evaluation methods assess the performance of tools based on the correctness of individual components of log statements. However, current metrics struggle to accurately reflect the quality of runtime logs in real execution environment. Even a slight shift of log statement can lead to the miss of essential runtime details, such as a several-lines shift from the ground truth or a mismatch in verbosity level. For instance, a minor difference in verbosity levels (e.g., debug vs. info) can cause critical information missed due to log level threshold settings in the source code. Therefore, we need to evaluate the quality of log statements in a real execution environment, with the goal of obtaining appropriate logs in specific scenarios, rather than merely focusing on the correctness of individual components or relying on statistical metrics to reflect the performance of logging tools.
Our Work. To address these challenges, we introduce AL-Bench: a comprehensive benchmark featuring a large-scale, diverse dataset and a novel approach for evaluating both the compilability and runtime logs produced by generated log statements. Our dataset comprises 42,224 instances collected from 10 popular, high-quality GitHub projects (GitHub, [n. d.]) spanning different domains with varying logging requirements, providing a robust foundation for evaluating automatic logging methods. Beyond the dataset, AL-Bench introduces a novel evaluation method called dynamic evaluation, which involves reintegrating generated log statements into real project code, followed by recompiling and executing them. This process allows for a realistic assessment of both compilability and runtime logs, highlighting major limitations in state-of-the-art tools: even the best tools fail to compile in 20.1% of cases, and the logs produced by generated log statements show only 0.213 cosine similarity to ground-truth logs. Through its rigorous and standardized evaluation approach, AL-Bench bridges the gap between real-world logging requirements and prior assessments, highlighting substantial opportunities to further advance the development of automatic logging tools.
This paper’s contributions are summarized as follows:
  1. (1)
    We collected a high-quality, diverse, and large-scale dataset comprising 42,224 instances from 10 popular, high-quality GitHub projects (GitHub, [n. d.]), spanning various domains with differing logging requirements.
  2. (2)
    We propose a novel dynamic evaluation approach that assesses generated log statements in real-world settings by reintegrating them into projects to evaluate their compilability and resulting runtime logs.
  3. (3)
    We conducted a comprehensive evaluation of popular automatic logging tools and revealed the key limitations based on the analysis of the evaluation results.
  4. (4)
    All the data and code for AL-Bench are publicly available  [^1] https://github.com/shuaijiumei/logging-benchmark-scripts, providing valuable resources for both developers and researchers to advance the field of automatic logging.
2. background and motivation
This section introduces the overview of the current research on automatic logging tools, followed by the shortcomings in the evaluation work, and explains the motivation for AL-Bench.
2.1. Log Statement Generation
Logging, the process of generating informative log messages with appropriate verbosity levels at strategically placed locations within code, has long been recognized as a critical challenge in software engineering (He et al., 2021; Chen and Jiang, 2019, 2021; Chen et al., 2021). Over the years, substantial research efforts have aimed to support developers in crafting more effective logging statements, which in turn enhance software maintenance and testing (Jia et al., 2018; Zhao et al., 2017; Zhu et al., 2015). Early studies in this domain often addressed isolated subproblems, typically operating under stringent assumptions that limit the applicability of their findings in real-world scenarios. For example, Li et al. (Li et al., 2021) proposed DeepLV to predict the appropriate logging level by taking surrounding code features into a neural network. Liu et al. (Liu et al., 2022) proposed Tell to further adapted flow graphs to help the suggestions of verbosity levels. Zhu et al. (Zhu et al., 2015) proposed LogAdvisor and Yao et al. (Yao et al., 2019) proposed Log4Perf to assist developers add new log statements in a specific position. Ding et al. proposed LoGenText (Ding et al., 2022) and LoGenText-Plus (Ding et al., 2023) to advise developers what should be logged, and Liu et al. (Liu et al., 2021) proposed tools for deciding which variables should be logged. However, none of them can generate a complete log statement.
Recently, Mastropaolo et al. (Mastropaolo et al., 2022) proposed the first end-to-end tool LANCE to generate complete log statements based on T5 (Raffel et al., 2019). Following LANCE, Xu et al. (Xu et al., 2024) proposed UniLog to adapt ICL and warm-up strategy to enhance the LLM ability for generating log statements. Xie et al. (Xie et al., 2023) proposed FastLog increase the generation time while keeping the accuracy, and Mastropaolo et al. (Mastropaolo et al., 2023) further proposed LEONID with a combination of Deep Learning (DL) and Information Retrieval (IR) achieving a better performance.
While end-to-end automatic logging tools have demonstrated promising results in their respective evaluations, our analysis reveals notable issues in both the datasets used and the evaluation methods employed.
Figure 2. The low-quality examples in the LANCE dataset follow the strictest data collection rules.
2.2. Limitations of Evaluation Methodology
2.2.1. Evaluation Data
Current evaluation datasets are ad-hoc, derived from splitting the entire collected data according to loose standards. More specifically, to ensure a sufficient amount of training data, previous studies had to adopt lenient rules for data collection. Additionally, some tools, to accommodate the limitations of their backbone models, further filter out data exceeding 512 tokens in length. These strategies compromise data quality, reducing the reliability and effectiveness of evaluation results. For example, LANCE adopted the strictest data collection rules among all tools, requiring a minimum of 500 commits, 10 contributors, and 10 stars, while also excluding forked repositories on GitHub (GitHub, [n. d.]). However, even in the LANCE dataset, low-quality data is pervasive. As shown in Figure 2, several patterns of low-quality logging practices persist. These patterns include the duplication of variables within a single log statement, creating unnecessary redundancy; the inclusion of empty strings or meaningless content, resulting in uninformative messages; excessive use of special characters or punctuation, making printed logs difficult to parse; and mismatched logging levels, where critical messages are assigned inappropriate severities, leading to misclassification (Adzic, 2006; Chen and Jiang, 2017). Using low-quality data as an evaluation dataset fails to produce reliable results. Such results do not align with the expectation that logging tools should generate high-quality log statements within the code. Furthermore, to accommodate the input length limitations of these tools, previous studies (Mastropaolo et al., 2022, 2023; Xu et al., 2024; Xie et al., 2023) have filtered out instances longer than 512 tokens. While this data-cleaning strategy reduces the complexity of evaluation data, it also introduces significant biases by excluding longer code snippets, which are common in real-world development environments. This approach simplifies the evaluation process but fails to capture the full scope of challenges that logging tools would face when dealing with complex and extended code bases. Consequently, the evaluation results may not accurately reflect the tools’ ability to handle the demands of real-world software development, where longer snippets and more intricate code structures are prevalent. Therefore, there is a pressing need for a public, large-scale, high-quality, and diverse benchmark dataset that can better represent real-world codebases and provide a standardized platform for evaluating automatic logging tools.
Figure 3. An example of the limitation of static evaluation, where only the quality of log statements is considered. It cannot assess the compilability of the code or quality of the printed runtime logs.
2.2.2. Evaluation Method
The current evaluation method focuses on assessing model performance by comparing the accuracy of each log statement component (i.e., position, verbosity level, message). We refer to this as static evaluation in this paper. Figure 3 provides an example of static evaluation to demonstrate how it works. In this example, only the verbosity level matches, so the matched level count will increase by one, and BLEU scores will be used to calculate the average score. Finally, the matched portion of each component and the average BLEU score will be combined to reflect the performance of the tools. These statistical metrics provide insights into evaluating the quality of predicted log statements; however, this approach still does not align with the goal of ensuring that the predicted log statements generate high-quality logs.
Figure 4. An example of Compilation Failure: Unreachable Statements
First, it cannot assess the compilability of predicted log statements. Compilability is a fundamental requirement for generating appropriate logs. This capability ensures that log statements could be applied into the source code without introducing compilation errors, avoiding wasting developers efforts to debug for the predicted log statements. However, the current evaluation methodology, which focuses solely on comparing the individual components of log statements, cannot identify predicted log statements that use undefined variables or contain incorrect code syntax which could introduce compilation errors. As shown in Figure 4, the predicted log statement is injected into an unreachable position. Applying static evaluation would fail to recognize that this prediction could cause compilation errors. Instead of classifying it as a completely incorrect prediction, the evaluation would count it as a level match instance and assign it a BLEU score of 55.03, contributing to the statistical value. This could introduce significant bias when evaluating the performance, appearing favorable in metrics but performing poorly in actual usage.
Second, it cannot assess the runtime logs generated by predicted log statements. The goal of writing log statements is to obtain the appropriate logs during software execution, so that the quality of log statement is ultimately decided by the quality of runtime logs. Although comparing the correctness of log statement components provides insights into assessing the quality of logs, it can introduce bias when using statistical metrics that focus solely on log statements to evaluate the performance of logging tools. A simple example is a mismatched verbosity level between the predicted log statement and the ground truth. In a case where only the verbosity level is mismatched while the other components are exactly matched, this instance would receive a high score in static evaluation. However, in a real-world scenario, for important error logs, if the predicted level is below the threshold, the logs will not be generated, regardless of how critical the message is. Conversely, for debug logs, if the predicted level exceeds the threshold, these logs could be printed in the production environment, leading to issues such as sensitive information leakage and increased storage costs. These two situations demonstrate that even a shift in verbosity level can turn a log statement into a complete bad case. A more complicated case is presented in Figure 5. Under static evaluation, this prediction with only shift in two lines in position, also could be rewarded highly in statistic metrics. However, it is completely incorrect: placing the log statement at line 2 risks a NullPointerException if the timer object is null when timer.get() is called. This would cause the program to crash during execution, meaning that not only would the logs fail to be generated, but the entire program could terminate unexpectedly. In real-world scenarios, this kind of error is critical and goes unnoticed in static evaluations. While static evaluation rewards this prediction for structural correctness, it fails to account for the actual behavior of the code when executed. Therefore, this example reinforces the need for more comprehensive evaluation methods that go beyond code-level metrics and consider the real-world implications of log statements.
In a conclusion, a log statement’s quality should be determined by its ability to avoid introducing extral errors and generate meaningful, contextually appropriate logs.
Figure 5. An example of the limitation of static evaluation. The prediction differs from the ground truth by a slight upward shift of two lines. It causes a NullPointerException when timer is null and fails to generate logs.
Insights: The motivating study underscores major limitations in current evaluation efforts, including the lack of a diverse, high-quality dataset and methods that do not meet real-world needs. To improve log statement evaluation, we should collect a high-quality, diverse dataset and use an execution-based method to directly evaluate the runtime logs generated by log statements.
3. AL-Bench
In this section, we introduce our benchmark, AL-Bench, which builds on insights from previous studies (Mastropaolo et al., 2022, 2023; Xu et al., 2024; Xie et al., 2023). AL-Bench consists of two key parts: static evaluation and dynamic evaluation, both of which we will introduce in detail. The static evaluation component includes a high-quality dataset with 42,224 instances and five metrics, based on the static evaluation method outlined in Section 2.2.2. The dynamic evaluation component introduces a separate dataset with 2,238 instances and four metrics, following the dynamic evaluation approach, which will be described later in this section.
3.1. Static Evaluation
3.1.1. Static Evaluation Method.
Static evaluation focuses on log statement components—position, verbosity, and message—and has been the primary method in prior studies (Mastropaolo et al., 2022, 2023; Xu et al., 2024; Xie et al., 2023), with details provided in Section 2.2.2.
3.1.2. Dataset Construction
As discussed in Section 2.2.1, the quality of the evaluation dataset is crucial for assessing the performance of tools. The strictest existing rules for dataset collection—500 commits, 10 contributors, and 10 stars—are inadequate to ensure the quality of the evaluation dataset. Drawing inspiration from the use of GitHub repository stars as an effective metric for identifying high-quality code datasets  (Jiang et al., 2024), and considering the unique characteristics of log statements, we propose three critical criteria for dataset selection: repositories must have at least 10,000 stars, 1,000 log statements, and 500 log-related issues to qualify for inclusion in our dataset. These standards guarantee well-maintained, widely used projects where developers prioritize high-quality log statements. Given the diverse requirements for log statements in different scenarios, we ultimately selected 10 projects based on domain diversity, industrial applications, and their relevance to prior logging studies (Chen and Jiang, 2017, 2019; Li et al., 2018, 2019; Ding et al., 2022, 2023; Li et al., 2020). As shown in Table 1, our final dataset includes projects with a total of 22,787 code snippets and 42,224 log statements, covering a wide range of logging needs and practices. The dataset spans multiple domains, including database management, task scheduling, distributed storage, messaging systems, and IoT platforms, ensuring diversity in logging scenarios. Each domain presents unique requirements for log statements.
For example, database management systems such as DBeaver and Doris prioritize minimizing the impact of logging on high performance. Task scheduling systems, including DolphinScheduler, rely on logging to trace task dependencies and monitor runtime statistics. Similarly, distributed systems like Hadoop and Zookeeper require robust logging practices to address challenges in distributed coordination, fault tolerance, and scalability, which differ from the requirements of other command logs. Messaging systems such as Kafka and Pulsar have adapted logging practices to trace message flows, ensure reliable message delivery, and debug asynchronous communication. Meanwhile, IoT platforms like ThingsBoard utilize logging to manage device connectivity, monitor data streams, and enable real-time system oversight. Identity and access management systems such as Keycloak prioritize protecting sensitive information in logs to prevent privacy breaches. By including diverse projects across these domains, the dataset ensures comprehensive coverage of different logging practices, making it suitable for benchmarking and analyzing log-related tasks under various operational contexts.
In addition to emphasizing data quality, we also addressed the potential risk of data contamination. Since all our data were extracted from public GitHub repositories, which may have been used for training pre-trained models, we implemented precautions to minimize this risk. Specifically, we collected the latest version of each project to reduce the likelihood of it being included in any model’s pre-training data. Furthermore, we wrapped the code snippets in a common class, named A, and standardized the formatting using Google-Java-Format (Google, [n. d.]). This approach altered the format of the code to prevent pre-trained models from recognizing the same information and structure. These strategies have been demonstrated as effective in recent studies (Xu et al., 2024; Tian et al., 2024) Although we adopted effective strategies, data contamination cannot be entirely avoided (Cao et al., 2024). However, our methods minimize this risk and have been proven effective in previous work, providing a solid foundation for analysis and evaluation. In the future, we plan to regularly update our dataset to ensure that the evaluation data remains current. After completing the necessary appeal actions, we finalized our static evaluation dataset.
Table 1. Details of AL-Bench datasets
Dataset Domain Code snippets with log statements Total log statements
Dbeaver Database Management 1,182 1,725
Dolphinscheduler Task Scheduling 898 1,918
Doris High Performance Database 1,686 2,965
Flink Data Processing 1,956 3,268
Hadoop Distributed Storage 8,751 16,002
Kafka Messaging Systems 1,706 3,421
Keycloak Identity and Access Management 533 1,004
Pulsar Messaging Systems 3,590 7,254
Thingsboard IoT Platform 1,601 2,731
Zookeeper Distributed Coordination 884 1,936
Total - 22,787 42,224
3.1.3. Metrics
Building on previous studies (Mastropaolo et al., 2022, 2023; Xu et al., 2024; Xie et al., 2023), We adopted five metrics for static evaluation. In addition to the previously established metrics, we propose two new ones: Dynamic Variable Accuracy and Static Text BLEU, designed to better reflect the real-world quality requirements of log statements.
Metric 1: Level Accuracy (LA): Level Accuracy focuses on log level values, which indicate the importance of the message and are a crucial part of log statements. Some common levels, like ”info,” refer to the normal information of runtime behavior. ”Warning” indicates potential problems that might not immediately cause a disruption but could lead to future issues if not resolved. ”Error” refers to runtime anomalies or issues that need to be addressed. Each level refers to different meanings so that for Level Accuracy we strictly compared the level of predictions and source code and using the exactly matched level number $L_{c}$ divided by the total number of log statements $N_{a}$ to get the value of this metric: $LA=\frac{L_{c}}{N_{a}}$.
Metric 2: Position Accuracy (PA): Position Accuracy focuses on the precise location of log statements within the source code, which is crucial for tracking and debugging software behavior. Correct placement of log statements helps in accurately tracing the execution flow and diagnosing issues. For Position Accuracy, we rigorously compare the predicted positions of log statements with their actual positions in the source code. This metric is calculated by taking the number of correctly positioned log statements $P_{c}$ and dividing it by the total number of log statements $N_{a}$ to obtain the accuracy value: $PA=\frac{P_{c}}{N_{a}}$.
Metric 3: Message Accuracy (MA): Message Accuracy evaluates how accurately the predicted log messages match the ground truth log messages, which is essential for providing meaningful and relevant information during runtime. The content of log messages helps developers understand the system’s behavior, and inaccuracies in message generation can lead to confusion or missed insights during debugging. For Message Accuracy, we compare the predicted log messages to the actual messages in the source code. This metric is calculated by determining the number of log messages that are fully identical to the ground truth $M_{c}$ and dividing it by the total number of log messages $N_{a}$, yielding the accuracy value: $MA=\frac{M_{c}}{N_{a}}$.
Metric 4: Dynamic Variable Accuracy (DVA): Dynamic Variable Accuracy focuses on the dynamic variable in log message, for example, (”The server is running, {}”, status), in this message, status is regarded as the dynamic variable in log message which might vary according to the runtime behaviors of software. And in the log message (”The server run on the ports, {}”, args.status ? localPort : remotePort), the whole expression (args.status ? localPort : remotePort) is reckoned as the dynamic variable which could decide the port information recorded in log files. We aim to use this metric to ensure that the dynamic information recorded in logs remains consistent. We extract the dynamic variables from both the source code and predictions and then compare them to record instances where they match exactly. This metric is calculated by taking the number of the exactly matched $DP_{c}$ and dividing it by the total number of log statements $N_{a}$ to obtain the accuracy value: $DVA=\frac{DP_{C}}{N_{a}}$.
Metric 5: Static Text BLEU (STB): Static Text BLEU focuses on the static part of the log message. Unlike the dynamic variable, the static part always records the same information in log files, which will not vary due to the runtime behavior of software. For example, in the log message (”The server is running, {}”, status), ”The server is running, {}” is regarded as the static part. Since this part primarily consists of natural language content, we use BLEU (Papineni et al., 2002) metric to evaluate the quality of the static text. In our implementation, we use the BLUE-DM variant (Chen and Cherry, 2014; Shi et al., 2021), i.e., the sentence-level BLEU without any smoothing method.
3.2. Dynamic Evaluation
Figure 6. The general workflow of dynamic evaluation. First, compile the project and run the unit test to obtain ground truth logs. Then, replace log statements with predictions, rerun the test to generate new logs, and finally analyze the results.
3.2.1. Dynamic Evaluation Method
Different from static evaluation, dynamic evaluation focuses on compiling the code and runtime-generated logs, addressing static evaluation’s inability to verify code compilability and runtime logs. To directly assess runtime logs, we generate them using unit tests, which are widely used in software development to verify code functionality in isolated scenarios. Unit tests are readily available in most projects and they are designed to test the functionality and behaviors of code when facing different situations. They offer a natural method for simulating realistic situations, allowing generating logs without the need for complex runtime environments. Figure 6 demonstrates the general workflow of dynamic evaluation. The process begins with compiling the source code and executing the unit tests to obtain logs from the original log statements. Using tools like Jacoco (Jacoco, [n. d.]) and SureFire (surefire, [n. d.]), we collect the logs and remove the log statements covered by unit tests in the source code. Next, we input the modified source code into an automatic logging tool to generate predicted log statements. Then we inserted the predicted log statements back into the source code, replacing the original log statements. After recompiling the modified code, we rerun the unit tests to capture the logs produced by the inserted predicted log statements. The whole process provides us with two sets of logs—those generated by the original log statements and those generated by the predicted log statements. Finally, we evaluate the effectiveness of the predicted log statements using two key metrics: Compilation Success Rate and Log File Similarity. Compilation Success Rate ensures that the predicted log statements do not introduce compilation errors, while Log File Similarity measures the similarity between the logs generated by the predicted log statements and those generated by the original log statements. In the following sections, we will detail how we built the dataset for dynamic evaluation and introduce the specific metrics used to measure the performance of the automatic logging tools.
3.2.2. Dataset Construction
To build the dynamic evaluation dataset, we begin by compiling the entire project to ensure all dependencies are resolved and the project is ready for execution. Next, we systematically identify all available unit tests within the project. For each unit test, we execute it individually while employing the Jacoco Plugin (Jacoco, [n. d.]) to trace code coverage, specifically identifying whether the unit test interacts with or covers any log statements in the codebase. Simultaneously, we use the SureFire Plugin (surefire, [n. d.]) to capture the logs generated during the execution of the unit tests.
By correlating Jacoco coverage data with SureFire logs, we can match specific code snippets containing log statements to the corresponding unit tests that cover them, along with the runtime logs they generate. This process enables us to construct a comprehensive dataset consisting of triples: the code snippet, the unit test that triggers it, and the recorded logs. These triples are critical for dynamic evaluation, as they provide the ground truth for assessing the quality and effectiveness of predicted log statements in actual runtime scenarios.
To construct a dynamic evaluation dataset, it is essential to select high-quality projects with comprehensive unit tests, as these tests provide a realistic simulation of diverse production environments. Each instance of dynamic evaluation requires recompiling the project and executing the test to collect logs, making the process highly time-consuming. To balance this intensive time requirement with the need for sufficient dataset diversity and quantity, we employed Hadoop as our dynamic evaluation platform. This approach allowed us to build a dataset of 2,238 instances that balances diversity with sufficient size. Additionally, we open-source the entire suite of tools used in this evaluation process, enabling researchers and organizations to easily deploy and customize their own dynamic evaluation datasets.
3.2.3. Metrics
In dynamic evaluation, we proposed four metrics to assess the performance of logging tools: Compilation Success Rate, Log Similarity, False Positive Log Generation Rate, and False Negative Log Generation Rate, which we will introduce below.
Metric 6: Compilation Success Rate (CSR): Compilation Success Rate measures the syntactic correctness of the predicted log statements. As discussed earlier, after replacing the original log statements with the predictions, we recompile the project. However, due to issues such as undefined variables in the predictions or missing/outdated dependencies in the project environment, not all predictions can be successfully compiled. We recorded the successfully compiled code snippet number as $C_{s}$ and all code snippets as $C_{a}$. The metric is then calculated as: $CSR=\frac{C_{s}}{C_{a}}$.
Metric 7: Log Similarity Metrics Group (Cosine Similarity, BLEU, ROUGE): This metric evaluates how closely logs generated by predicted statements match those produced by ground truth statements. To eliminate unnecessary differences, we remove log headers (e.g., timestamps), retaining only log content. For a comprehensive assessment, we apply multiple similarity measures, including Cosine Similarity (Salton et al., 1975), BLEU (Papineni et al., 2002), and ROUGE (Lin, 2004). Cosine Similarity, commonly used in text analysis, calculates the cosine of the angle between two TF-IDF (Jones, 2021) vectors, yielding 1 for identical vectors and 0 for orthogonal ones. Using TF-IDF, we down-weight frequent terms, emphasizing distinctive content in logs. This method effectively captures the similarity between meaningful log content, filtering out redundant information for a more accurate relevance measure. ROUGE, on the other hand, focuses on recall by comparing n-grams between the predicted and reference logs. It evaluates how much of the reference content is preserved in the prediction. The most commonly used variant is ROUGE-N, which calculates the overlap of n-grams between two texts.
Metric 8: False Positive Log Generation Rate (FPLR): This metric measures the proportion of predicted log statements that generate logs during unit test execution when the ground truth log statements would not have produced any logs. It helps assess whether the predicted log statements introduce unnecessary or redundant logs in scenarios where no log should be generated. The number of false positive instances is recorded as $FP$, and the total number of predictions is $P$. he metric is calculated as: $FRLR=\frac{FP}{P}$.
Metric 9: False Negative Log Generation Rate (FNLR): This metric evaluates the proportion of predicted log statements that fail to generate logs during unit test execution when the ground truth log statements should have produced logs. It highlights instances where the predicted logs miss important events or information. The number of false negative instances is recorded as $FN$, and the total number of predictions is $P$. The metric is calculated as: $FNLR=\frac{FN}{P}$.
4. EXPERIMENTS
In this section, we use AL-Bench to evaluate existing end-to-end automatic logging tools and analyze the evaluation results to find insights to guide the following work. From the perspective of empirical software engineering, we set three research questions:
  • RQ1: How well are logging tools at predicting log statements in static evaluation?
  • RQ2: How well can the predicted log statements be compiled successfully?
  • RQ3: How well do the generated log statements print logs in dynamic evaluation?
Specifically, RQ1 examines how well logging tools predict log statements using the AL-Bench static evaluation dataset and corresponding static evaluation method. This research question aims to evaluate the accuracy and effectiveness of the tools in generating log statements that closely resemble the ground truth. RQ2 focuses on assessing the compilability of the predicted log statements, determining whether these predictions can be seamlessly integrated into the code without causing compilation errors. Finally, RQ3 evaluates the runtime logs produced by the generated log statements. We examine not only the generation of inappropriate logs but also the similarity between the predicted logs and the expected logs in cases where logs are correctly generated. We begin by introducing the automatic logging tools selected for evaluation. We will provide a detailed analysis of these tools, focusing on their performance across the three research questions. Each tool is assessed for its ability to predict accurate log statements, compile them successfully, and generate meaningful logs. This comprehensive evaluation allows us to draw insights into the strengths and limitations of current approaches to automatic logging.
4.1. Evaluating Automatic Logging Tools
Automatic logging is a hot topic, leading to the development of many tools for determining specific parts of log statements and end-to-end automatic logging tools in recent years. In this paper, we focus on end-to-end logging tools. We reached out to the authors of popular end-to-end automatic logging tools for assistance in rebuilding these tools. However, due to security policies, SCLogger (Li et al., 2024) is still under construction. We ultimately selected four methods for evaluation: LANCE (Mastropaolo et al., 2022), LEONID (Mastropaolo et al., 2023), FastLog (Xie et al., 2023), UniLog (Xu et al., 2024). We detailed the method in the following.
LANCE: $LANCE$ (Mastropaolo et al., 2022) is the first model designed to generate and insert complete log statements in code. It takes a method requiring a log statement and outputs a meaningful log message with an appropriate logging level in the correct position. Built on the Text-To-Text Transfer Transformer model, $LANCE$ is trained specifically for injecting proper logging statements.
LEONID: $LEONID$ (Mastropaolo et al., 2023) is the updated version of $LANCE$. With a combination of DL and Information Retrieval (IR), $LEONID$ achieved a better performance. $LEONID$ provided two versions, $LEONID_{S}$ is for single log statement generation, and $LEONID_{M}$ is for multiple log statements generation. Since $LEONID_{M}$ can generate more than one log statement at a time, it introduces ambiguity in determining the correct correspondence between the generated and expected log statements when more than one log statements are generated by static evaluation (Mastropaolo et al., 2023). Therefore, we only applied $LEONID_{S}$ for static evaluation.
UniLog: $UniLog$ (Xu et al., 2024) is the first attempt to adapt Warm-up and In-context-learning strategy to enhance the model’s ability to generate log statements. Due to limitations in assessing the original UniLog, we reproduced it using two backbone models: CodeLlama-7B (Rozière et al., 2023) and DeepSeek-V3 (DeepSeek-AI et al., 2024). We applied the warmup process exclusively to the CodeLlama backbone model, while employing the ICL strategy to construct prompts for both models. The data are sourced from LANCE (Mastropaolo et al., 2022) to warm up and generate ICL content. The effectiveness of In-Context Learning often depends on whether the examples are in-distribution or out-of-distribution relative to the evaluation data. Since LANCE’s data distribution differs from our evaluation data, this may affect UniLog’s performance. We will use $UniLog_{{cl}}$ to represent the version based on CodeLlama-7B, $UniLog_{{ds}}$ to represent the version based on DeepSeek-V3.
FastLog: FastLog (Xie et al., 2023) defines the logging task in two steps: finding the position and generating and inserting a complete log statement into the source code. This approach avoids rewriting the source code, a key limitation of $LANCE$. They utilized PLBART (Ahmad et al., 2021) as the base model to fine-tune two separate models: one for predicting insertion position, the other for generating log statements. With the heuristic rule, log statements only appear after certain special characters, FastLog enhances efficiency while maintaining accuracy in generating log statements.
Table 2. Static evaluation on the complete logging task was conducted. The values in the ‘Original’ lines indicate the reported results from the respective methods’ own evaluations. The best performance evaluated by AL-Bench across different metrics is highlighted.
Method PA LA MA DVA STB (BLEU)
Now Original Now Original Now Original
FastLog 57.54 $58.84$ 62.72 $59.75$ 6.90 $4.52$ 17.66 20.20
$UniLog_{{ds}}$ $36.31$ $76.90$ $36.31$ $72.30$ $5.19$ $22.40$ 15.73 11.60
$UniLog_{{cl}}$ $23.29$ $76.90$ $23.29$ $72.30$ $2.31$ $22.40$ 16.08 8.43
LANCE $34.67$ $65.40$ $34.67$ $66.24$ $2.99$ $16.90$ 14.42 6.52
$LEONID_{S}$ $15.22$ $76.45$ $15.22$ $73.53$ $1.38$ $31.55$ 5.35 2.03
4.2. RQ1: How well are logging tools at predicting log statements in static evaluation?
Figure 7. Performance of logging tools among different projects. The performance of each tool varies considerably across different projects and the trends of all methods across all projects generally remain consistent.
We evaluated four automatic logging tools using the AL-Bench static evaluation dataset and corresponding static evaluation methods, ensuring a fair comparison of their performance.
Following previous studies (Mastropaolo et al., 2022; Xu et al., 2024; Xie et al., 2023; Mastropaolo et al., 2023), we first processed our dataset to create evaluation pairs $<M_{s},M_{t}>$ with $M_{s}$ representing the input provided to the model (i.e., $M_{s}$ with one removed log statement) and $M_{t}$ being the expected output (i.e., $M_{t}$ is the removed log statement). After processing, we got 42,224 instances. The evaluation results are presented in Table 2. The metrics for Position Accuracy (PA), Level Accuracy (LA), and Message Accuracy (MA) are adapted from the evaluation frameworks used by the selected logging tools. Therefore, we label the values of these metrics as reported in their respective evaluation studies. The results demonstrate significant performance differences among the four logging tools.
FastLog consistently outperforms the other tools across all metrics, but it still faces challenges in specific areas. For instance, it scores only 17.66 in Dynamic Variable Accuracy (DVA) and 20.44 in Static Text BLEU (STB), suggesting that even the best tool struggles to generate meaningful natural language descriptions and accurately select dynamic variables for logging. Despite these shortcomings, FastLog shows only minor deviations from previous studies, maintaining stable performance across all metrics. In contrast, the other three tools, particularly LEONID, experience a significant decline in accuracy across all evaluated metrics.
As illustrated in Figure 7, the performance of each tool varies considerably across different projects. This fluctuation suggests that the reliability and stability of logging tools remain inconsistent when applied to diverse scenarios. The performance fluctuations of logging tools across different projects are due to variations in project complexity and the diverse logging requirements. Additionally, limited training data diversity may hinder the tools’ ability to perform consistently across unfamiliar scenarios. These factors collectively contribute to the observed instability in logging tool performance across different projects. As demonstrated in our results, the variability in tool performance across different projects underscores the necessity of diverse evaluation data to capture the full spectrum of a tool’s capabilities and limitations.
To investigate the causes of performance differences compared to previous studies, we observed that earlier evaluation datasets were filtered to exclude instances longer than 512 tokens. To assess the impact of this filtering, we divided our dataset into two groups: instances with shorter than 512 tokens and those with longer than 512 tokens. This division enables us to determine whether instance length contributes to the performance discrepancies observed among the logging tools, providing further insight into how different data characteristics influence tool effectiveness.
Table 3. The performance of tools when facing different length input data. Long means data longer than 512 tokens, Short means data shorter than 512 tokens, and $\Delta$ means the difference in logging tools performance when facing the Short data and the Long data.
Method Dataset PA LA MA DVA STB
FastLog Long 43.36 64.92 7.02 11.98 20.97
Short 61.79 62.33 6.88 19.36 19.98
$\Delta$ $\downarrow 18.43$ $\uparrow 2.59$ $\uparrow 0.14$ $\downarrow 7.38$ $\uparrow 0.99$
$UniLog_{d}{}_{s}$ Long 22.28 66.66 8.82 15.56 17.05
Short 40.51 59.52 4.11 15.78 9.97
$\Delta$ $\downarrow 18.23$ $\uparrow 7.14$ $\uparrow 4.11$ $\downarrow 0.22$ $\uparrow 7.08$
$UniLog_{c}{}_{l}$ Long 8.70 51.62 2.84 12.91 8.20
Short 27.67 51.78 2.62 17.02 8.80
$\Delta$ $\downarrow 18.97$ $\downarrow 0.16$ $\uparrow 0.22$ $\downarrow 4.12$ $\downarrow 0.60$
LANCE Long 0 0 0 0 0
Short 45.05 47.61 3.89 18.74 8.12
$\Delta$
$LEONID_{S}$ Long 0 0 0 0 0
Short 19.78 21.46 1.79 6.95 2.64
$\Delta$
After dividing the dataset, we obtained two groups: a longer dataset containing 9,732 instances and a shorter dataset with 32,492 instances. The evaluation results in Table 3 show a clear difference in tool performance based on the length of the instances. Notably, LANCE and LEONID struggled with instances longer than 512 tokens, failing to generate syntax-correct code and, in some cases, producing incomplete code. This explains why their scores for these cases are reported as zero, highlighting the input length limitations of both tools and their inability to handle longer, more complex instances effectively. $UniLog_{cl}$ and FastLog show a considerable drop in PA and DVA when handling longer data, indicating that they struggle to predict log positions and select dynamic variables in more complex instances. $UniLog_{ds}$also shows a dramatic drop in position accuracy, but only a slight decrease in the ability to decide dynamic variables for recording. However, with more sufficient information and a more powerful backbone model, it demonstrates significantly better performance in other metrics. This indicates that with a more powerful backbone, UniLog can better understand the context with longer inputs. Although $UniLog_{d}s$ is equipped with a more powerful backbone, it still struggles to choose the appropriate position for the log statement. As widely demonstrated that the LLMs are not good at counting numbers (Ahn et al., 2024), we might change the output format by directly adding log statement into the source code rather than generate the exactly line number. Furthermore, by simply analyzing the control flow graph of the code, we might be able to exclude positions where logging is not feasible to leave less choices.
Answer to RQ1. FastLog performs best in AL-Bench ’s static evaluation but still faces challenges in generating meaningful descriptions and selecting key dynamic variables. All tools struggle to maintain stable performance under varying logging requirements and complex data, especially in determining log positions. Since LLMs struggle with counting, it might be better to avoid outputting exact line numbers and instead add tags in the source code. Additionally, analyzing the control graph to exclude impossible positions might be useful to enhance log position determination.
Table 4. Compilation Failure Rates across methods
Methods Compilation Success Rates
FastLog 79.9%
$UniLog_{{cl}}$ 70.3%
$UniLog_{{ds}}$ 60.2%
LANCE 49.4%
$LEONID_{M}$ 25.0%
$LEONID_{S}$ 16.4%
Table 5. Compilation Failure Reasons Analysis
Failed Reason Failed Number
Using Wrong Logging Name 56
Using Undefined Method 21
Using Undefined Variables 15
Incompatible Types 3
Unreachable Statements 1
Other 4
Total 100
4.3. RQ2: How well can the predicted log statements be compiled successfully?
To evaluate tools’ ability to generate compilable log statements, we replace existing logs with predicted ones in the dynamic evaluation dataset and recompile the project to check for successful compilation. As shown in Table 5, the best-performing tool, FastLog, achieves a 79.9% Compilation Success Rate, followed by $UniLog_{c}l$ at 70.3%, $UniLog_{d}s$ at 60.2%, LANCE at 49.4%, $LEONID_{M}$ and $LEONID_{S}$ at 25.0% and 16.4%, respectively. The results reveal a key limitation of LANCE and LEONID that they regenerate the entire code snippet, which increases the risk of unintended code changes and can potentially lead to compilation errors. In contrast, FastLog and UniLog focus solely on generating the new log statement, minimizing the risk of errors by limiting codebase modifications. Although FastLog is one of the best tools available, it still leads to significant instances of compile failure. When log statements fail to compile, they can prevent the system from functioning as intended and complicate debugging or error tracking. To ensure the reliability of automatic logging tools in real-world environments, it’s crucial that the log statements they generate integrate smoothly into the existing codebase.
Given the importance of reliable compilation, we conducted a manual review of the compilation failures to identify the specific causes. This analysis will provide further insight into the key weaknesses of these tools and potential areas for improvement. We randomly selected 100 failed instances from the best tool, FastLog, and the first two authors cross-checked the causes. The analysis results are presented in Table 5. The most common failure, occurring in 56 instances, is due to Using the wrong Logging Name (i.e., Logger, LOG), indicating that incorrect or non-existent log functions are being invoked. Using Undefined Methods accounts for 21 failures, followed by Using Undefined Variables with 15 failures. Less frequent issues include Incompatible Types (3), and Unreachable Statements (1). Others mean generating the wrong syntax code, which we will not analyze.
The majority of failures (totaling 92 instances) involve undefined references to methods (21 instances), variables (15 instances), or logging names (56 instances). These undefined reference failures are primarily due to the limited context provided by existing function-level logging tools, lacking critical details on valid variables, methods, libraries, and packages relevant to the target function. Current tools are designed for function-level input, highlighting the need for logging tools to integrate better context awareness and validation checks to ensure compatibility with the existing codebase. The less frequent errors, such as incompatible types and unreachable statements, also indicate challenges in generating log statements that integrate well with the surrounding code logic. The issue of Unreachable Statements is notable because it points to a fundamental weakness in the current logging tools—specifically, their lack of understanding of the code’s control flow.
Answer to RQ2. FastLog achieved the best performance in generating compilable log statements, yet over 20% of the generated log statements still failed to be compiled. According to our analysis of compilation failure reasons, these failures primarily stem from a lack of critical contexts corresponding to the target function, e.g., valid variables, methods, libraries, packages, execution paths, and type information. To improve the reliability of automatic logging tools, it is crucial that they incorporate mechanisms to gather and utilize this additional context during the log statement generation process.
4.4. RQ3: How well do the generated log statements print logs in dynamic evaluation?
To answer RQ3, we used the dynamic evaluation dataset including code snippets paired with corresponding unit tests to generate logs. For each instance, we replaced the existing log statements in the source code with the predicted log statements from the automatic logging tools. The unit tests were then executed to observe the logs generated by the modified code.
It is important to highlight that, aside from $LEONID_{M}$, all other tools operate under the strong assumption that the given code snippet requires exactly one log statement. To highlight this inappropriate assumption, we only allow the tools one chance to predict the log statement, even in cases where multiple log statements might be needed. This experimental setup closely mirrors real-world conditions and allows for a more thorough evaluation of the tool’s ability to handle dynamic and varied logging requirements, moving beyond the oversimplified assumption that each snippet requires only one log statement. ultimately limiting tool performance in practical applications.
Table 6. The Semantic Similarity of Logs Printed by the Predicted Log Statements Across Methods.
Method Cosine Similarity BLEU BLEU-1 BLEU-4 ROUGE-1 ROUGE-L
$FastLog$ 0.213 16.172 24.829 15.963 24.427 23.841
$UniLog_{{cl}}$ 0.174 13.625 19.015 13.423 19.167 18.749
$UniLog_{{ds}}$ 0.130 11.607 14.674 11.556 14.002 13.869
$LANCE$ 0.099 8.201 11.342 7.971 11.235 11.033
$LEONID_{M}$ 0.072 5.466 7.946 5.303 8.095 7.944
$LEONID_{S}$ 0.044 3.220 4.908 2.955 5.099 5.012
Table 6 compares the semantic similarity between logs from source and predicted log statements. The results indicate that logs from the predicted statements significantly deviate from the ground truth, with consistently low scores across metrics like Cosine Similarity, BLEU, and ROUGE. Specifically, low similarity scores indicate that generated logs frequently fall short of matching expected outputs. Limiting each code snippet to a single log statement often sharply compromises log similarity, especially when multiple statements are needed. This limitation is particularly clear when comparing $LEONID_{S}$ and $LEONID_{M}$. Although $LEONID$ overall performs poorly, $LEONID_{M}$ stands out as the only tool capable of generating multiple log statements, which enables it to outperform $LEONID_{S}$ in similarity scores. This difference underscores the importance of tools being able to determine the appropriate number of log statements for accurate and effective logging, rather than assuming a single statement suffices.
Table 7. False Positive and False Negative Logging Rates across Methods. FPLG means the predicted log statement record logs when it should not, and FNLG means the predicted log statement does not record logs when it should.
Method FPLG Rate FNLG Rate
FastLog 9.28% 18.28%
$UniLog_{{cl}}$ 6.52% 30.59%
$UniLog_{{ds}}$ 3.21% 22.88%
LANCE 5.71% 19.29%
$LEONID_{S}$ 8.15% 8.69%
$LEONID_{M}$ 7.32% 11.6%
Table 8. FPLG and FNLG Reason Analysis. For FNLG, the major reason is beyond the execution path, and for FPLG, the major reason is lower verbosity level.
Situation Reason Number
FNLG Beyond Execution Path 35
Lower Verbosity Level 24
Wrong Code Format 4
FPLG Higher Verbosity Level 30
Beyond Execution Path 3
Total - 100
To understand the low semantic similarity scores, we examined the log generation process and found many instances where predicted logs were either redundant or missed key information present in the original logs. We quantify this issue using two metrics: False Positive Log Generation (FPLG) and False Negative Log Generation (FNLG), as reported in Table 8. For example, FastLog reports a 9.2% False Positive Log Rate (FPLR), meaning logs record redundant information in 9.2% of cases, and an 18.28% False Negative Log Rate (FNLR), indicating expected information in logs were missing in 18.28% of cases. We sampled 100 examples from FastLog, the state-of-the-art (SOTA) model, and manually analyzed the reasons for mismatches with the original logs. The results are presented in Table 8. We found that the primary reasons for failures in FNLG and FPLG differ significantly. For FNLG, the most common issue was the predicted log statements being beyond the execution path (35 cases), followed by lower verbosity levels (24 cases), and a smaller number caused by wrong code format (4 cases). In contrast, for FPLG, the main problem was higher verbosity levels (30 cases), with a few instances of log statements being beyond the execution path (3 cases). Overall, verbosity mismatches and execution path discrepancies were the dominant contributors, highlighting challenges in aligning predicted logs with actual logging requirements. The factors leading to FPLG and FNLG underscore a critical issue: while static metrics offer valuable insights into the quality of generated log statements, the actual logs are shaped by numerous contextual factors. Without a thorough understanding of the execution context, it is not possible to comprehensively evaluate the quality of log statements. Even minor discrepancies can cause significant deviations between generated logs and source logs. For instance, while the predicted log statement may capture the key information required to reflect system behavior, its effectiveness can be compromised if it is not positioned along a critical execution path or if its verbosity level is mismatched. In such cases, the log statement may fail to record essential information when key events occur. This issue majorly arises from the tools’ lack of awareness of verbosity thresholds and the control graph of the code, which limits their ability to adjust verbosity and determine appropriate log positions based on the context or execution requirements. These two limitations highlight that current logging tools lack the adaptability and context-awareness needed for effective real-world application.
Answer to RQ3. The best predicted log statements by FastLog achieve only 0.231 cosine similarity with the original logs. Many predictions record redundant information, while others miss key details. The missing key information is primarily due to the prediction being placed beyond the execution path during important events, while setting higher verbosity level in the log statements leads to redundancy. This result highlights that automatic logging tools still have significant room for improvement.
5. Threats to Vadility
  • Construct Validity: We use BLEU to assess the quality of the generated log messages. Although text similarity metrics may not fully capture the quality of the generated text (Gros et al., 2020; Roy et al., 2021), we follow previous works (Xie et al., 2023; Xu et al., 2024; He et al., 2018a; Ding et al., 2023) to use BLEU as widely-accepted quality measures for the generated log messages. We also adapt Cosine Similarity based on TF-IDF to evaluate the quality of log content. Since log content, excluding the log header, is primarily composed of natural language, cosine similarity effectively captures the semantic meaning (Salton et al., 1975).
  • Internal Validity: The second threat to validity concerns reproducing the baseline. To minimize inconsistencies, we adapted the released models from previous work (Xie et al., 2023; Mastropaolo et al., 2022, 2023) and sought guidance from the authors of the closed-source model UniLog (Xu et al., 2024) to reproduce the tool under supervision.
  • External Validity: In this work we evaluate the tools with 10 widely recognized Java projects in different domains. The results of each tools are examined in different requirements of logging practices. The novel evaluation method we propose is not limited to Java projects, we will explore its effects in other programming languages in the future.
6. Conclusion
This paper introduces AL-Bench, designed specifically for automatic logging tools. It includes a high-quality dataset and a novel dynamic evaluation method focused on runtime logs, addressing key limitations of prior studies and bridging the gap between real-world requirements and existing evaluation frameworks. This dynamic evaluation method assesses both the compilability of predicted log statements and their effectiveness in generating runtime logs. Using AL-Bench, we evaluate popular end-to-end automatic logging tools and find that generated log statements fail to compile in 20.1% to 83.6% of cases. Even the best predictions achieve only 0.213 cosine similarity between generated and ground-truth runtime logs. These results show that automatic logging still has a long way to go.
7. DATA AVAILABILITY
All the code and data used in our study are publicly available on https://github.com/shuaijiumei/logging-benchmark-scripts.

### v2

@@
 Logging, the practice of inserting log statements into source code, is critical for improving software reliability.
-Recently, language model-based techniques have been developed to automate log generation based on input code.
-Although these methods demonstrate promising results in isolated evaluations, their effectiveness diminishes when applied to ad-hoc low-quality data and code similarity-based evaluation methods.
-We consider a comprehensive evaluation benchmark should include (1) a high-quality, diverse, and large-scale dataset, (2) an assessment of the compilability of the code with inserted log statements, and (3) a runtime log-oriented evaluation method.
-To this end, this paper introduces AL-Bench, a comprehensive benchmark designed specifically for automatic logging tools.
+Recently, language model-based techniques have been developed to automate log statement generation based on input code.
+These tools show promising results in their own evaluation.
+However, current evaluation practices in log statement generation face significant challenges.
+The lack of a unified, large-scale dataset forces studies to rely on ad-hoc data, hindering consistency and reproducibility.
+Additionally, assessments based solely on metrics like code similarity fail to reflect real-world effectiveness.
+These limitations underscore the need for a comprehensive public benchmark to standardize evaluation.
+This paper introduces AL-Bench, a comprehensive benchmark designed specifically for automatic logging tools.
 AL-Bench includes a high-quality, diverse dataset collected from 10 widely recognized projects with varying logging requirements and introduces a novel dynamic evaluation approach.
-Different from the evaluation in existing logging papers, AL-Bench assesses both the compilability of the code with inserted log statements and the quality of the logs generated by them during runtime, which we believe can better reflect the effectiveness of logging techniques in practice.
-AL-Bench reveals significant limitations in the state-of-the-art tools The codes with log statements generated by the state-of-the-art tools fail to compile in 20.1%-83.6% cases.
-In addition, even the best-performing tool did not achieve high similarity between the runtime logs produced by the generated log statements and the ground-truth log statements, demonstrating a 0.213 cosine similarity.
+Different from the existing evaluations that focus only on components of log statements like code similarity, AL-Bench assesses both the compilability of the code with inserted log statements and the effectiveness of the logs generated by them during runtime, which we believe can better reflect the effectiveness of logging techniques in practice.
+AL-Bench reveals significant limitations in the state-of-the-art tools.
+The codes with log statements generated by the state-of-the-art tools fail to compile in 20.1%-83.6% cases.
+In addition, even the best-performing tool only achieves 0.213 cosine similarity between the runtime logs produced by the generated log statements and the ground-truth log statements.
 The results reveal substantial opportunities to further enhance the development of automatic logging tools.
@@
 As illustrated in Figure 1, automatic logging typically includes three steps: (1) determining the position, (2) deciding the verbosity level, and (3) specifying the message to be recorded.
-Leveraging the advanced text generation capabilities of large language models (LLMs) (Floridi and Chiriatti, 2020; Raffel et al., 2020), LANCE is the first end-to-end tool to integrate positioning, level selection, and message generation.
+Leveraging the advanced text generation capabilities of large language models (LLMs) (Floridi and Chiriatti, 2020; Raffel et al., 2020), LANCE (Mastropaolo et al., 2022) is the first end-to-end tool to integrate positioning, level selection, and message generation.
 UniLog (Xu et al., 2024) employed a warm-up and in-context learning (ICL) strategy to enhance performance.
@@
 To relieve developers from the heavy effort required to design and maintain log statements (Chen and Jiang, 2019, 2017), the basic requirement is to ensure that our tool can be seamlessly integrated into the DevOps process without introducing additional errors that require extra debugging effort from developers.
-Current evaluation methods merely focus on whether each component of log statements (i.e., position, verbosity level, message) matches the ground truth but cannot assess whether the generated log statements might introduce compilation errors such as wrong code format or using undefined variables.
+Current evaluation methods merely focus on whether each component of log statements (i.e., position, verbosity level, message) matches the ground truth but cannot assess whether the generated log statements might introduce compilation errors such as wrong code format or use undefined variables.
 Evaluating the compilability of predicted log statements reflects the effectiveness and reliability of logging tools, which are essential for practical use.
@@
 Current evaluation methods assess the performance of tools based on the correctness of individual components of log statements.
-However, current metrics struggle to accurately reflect the quality of runtime logs in real execution environment.
-Even a slight shift of log statement can lead to the miss of essential runtime details, such as a several-lines shift from the ground truth or a mismatch in verbosity level.
+However, current metrics struggle to accurately reflect the quality of runtime logs in the real execution environment.
+Even a slight shift of log statement can lead to the miss of essential runtime details, such as a several-line shift from the ground truth or a mismatch in verbosity level.
 For instance, a minor difference in verbosity levels (e.g., debug vs.
@@
   1. (1)
-    We collected a high-quality, diverse, and large-scale dataset comprising 42,224 instances from 10 popular, high-quality GitHub projects (GitHub, [n.
-d.]), spanning various domains with differing logging requirements.
+    We collected a high-quality, diverse, and large-scale dataset comprising 42,224 instances from 10 popular, high-quality GitHub (GitHub, [n.
+d.]) projects, spanning various domains with differing logging requirements.
   2. (2)
@@
 (Zhu et al., 2015) proposed LogAdvisor and Yao et al.
-(Yao et al., 2019) proposed Log4Perf to assist developers add new log statements in a specific position.
+(Yao et al., 2019) proposed Log4Perf to assist developers to add new log statements in a specific position.
 Ding et al.
@@
 Compilability is a fundamental requirement for generating appropriate logs.
-This capability ensures that log statements could be applied into the source code without introducing compilation errors, avoiding wasting developers efforts to debug for the predicted log statements.
+This capability ensures that log statements can be applied to the source code without introducing compilation errors, avoiding wasting developers’ efforts to debug for the predicted log statements.
 However, the current evaluation methodology, which focuses solely on comparing the individual components of log statements, cannot identify predicted log statements that use undefined variables or contain incorrect code syntax which could introduce compilation errors.
@@
 Second, it cannot assess the runtime logs generated by predicted log statements.
-The goal of writing log statements is to obtain the appropriate logs during software execution, so that the quality of log statement is ultimately decided by the quality of runtime logs.
+The goal of writing log statements is to obtain the appropriate logs during software execution so that the quality of log statement is ultimately decided by the quality of runtime logs.
 Although comparing the correctness of log statement components provides insights into assessing the quality of logs, it can introduce bias when using statistical metrics that focus solely on log statements to evaluate the performance of logging tools.
@@
 Conversely, for debug logs, if the predicted level exceeds the threshold, these logs could be printed in the production environment, leading to issues such as sensitive information leakage and increased storage costs.
-These two situations demonstrate that even a shift in verbosity level can turn a log statement into a complete bad case.
+These two situations demonstrate that even a shift in verbosity level can turn a log statement into a completely bad case.
 A more complicated case is presented in Figure 5.
-Under static evaluation, this prediction with only shift in two lines in position, also could be rewarded highly in statistic metrics.
+Under static evaluation, this prediction with only a shift in two lines in position, also could be rewarded highly in statistic metrics.
 However, it is completely incorrect: placing the log statement at line 2 risks a NullPointerException if the timer object is null when timer.get() is called.
@@
 Therefore, this example reinforces the need for more comprehensive evaluation methods that go beyond code-level metrics and consider the real-world implications of log statements.
-In a conclusion, a log statement’s quality should be determined by its ability to avoid introducing extral errors and generate meaningful, contextually appropriate logs.
+In conclusion, a log statement’s quality should be determined by its ability to avoid introducing extra errors and generate meaningful, contextually appropriate logs.
 Figure 5.
@@
 Specifically, we collected the latest version of each project to reduce the likelihood of it being included in any model’s pre-training data.
-Furthermore, we wrapped the code snippets in a common class, named A, and standardized the formatting using Google-Java-Format (Google, [n.
+Furthermore, we wrapped the code snippets in a common class and standardized the formatting using Google-Java-Format (Google, [n.
 d.]).
 This approach altered the format of the code to prevent pre-trained models from recognizing the same information and structure.
-These strategies have been demonstrated as effective in recent studies (Xu et al., 2024; Tian et al., 2024) Although we adopted effective strategies, data contamination cannot be entirely avoided (Cao et al., 2024).
+These strategies have been demonstrated as effective in recent studies (Xu et al., 2024; Tian et al., 2024).
+Although we adopted effective strategies, data contamination cannot be entirely avoided (Cao et al., 2024).
 However, our methods minimize this risk and have been proven effective in previous work, providing a solid foundation for analysis and evaluation.
@@
 After completing the necessary appeal actions, we finalized our static evaluation dataset.
-Table 1. Details of AL-Bench datasets
+Table 1. Details of AL-Bench dataset.
 Dataset Domain Code snippets with log statements Total log statements
@@
 Some common levels, like ”info,” refer to the normal information of runtime behavior.
-”Warning” indicates potential problems that might not immediately cause a disruption but could lead to future issues if not resolved.
+”Warning” indicates potential problems that might not immediately cause disruption but could lead to future issues if not resolved.
 ”Error” refers to runtime anomalies or issues that need to be addressed.
-Each level refers to different meanings so that for Level Accuracy we strictly compared the level of predictions and source code and using the exactly matched level number $L_{c}$ divided by the total number of log statements $N_{a}$ to get the value of this metric: $LA=\frac{L_{c}}{N_{a}}$.
+Each level refers to different meanings so for Level Accuracy we strictly compared the level of predictions and source code and used the exactly matched level number $L_{c}$ divided by the total number of log statements $N_{a}$ to get the value of this metric: $LA=\frac{L_{c}}{N_{a}}$.
 Metric 2: Position Accuracy (PA): Position Accuracy focuses on the precise location of log statements within the source code, which is crucial for tracking and debugging software behavior.
@@
 The number of false positive instances is recorded as $FP$, and the total number of predictions is $P$.
-he metric is calculated as: $FRLR=\frac{FP}{P}$.
+the metric is calculated as: $FRLR=\frac{FP}{P}$.
 Metric 9: False Negative Log Generation Rate (FNLR): This metric evaluates the proportion of predicted log statements that fail to generate logs during unit test execution when the ground truth log statements should have produced logs.
@@
 This comprehensive evaluation allows us to draw insights into the strengths and limitations of current approaches to automatic logging.
-4.1. Evaluating Automatic Logging Tools
+4.1. Automatic Logging Tools
 Automatic logging is a hot topic, leading to the development of many tools for determining specific parts of log statements and end-to-end automatic logging tools in recent years.
@@
 We reached out to the authors of popular end-to-end automatic logging tools for assistance in rebuilding these tools.
-However, due to security policies, SCLogger (Li et al., 2024) is still under construction.
-We ultimately selected four methods for evaluation: LANCE (Mastropaolo et al., 2022), LEONID (Mastropaolo et al., 2023), FastLog (Xie et al., 2023), UniLog (Xu et al., 2024).
+Because SCLogger (Li et al., 2024) is still under construction, we ultimately selected four methods for evaluation: LANCE (Mastropaolo et al., 2022), LEONID (Mastropaolo et al., 2023), FastLog (Xie et al., 2023), UniLog (Xu et al., 2024).
 We detailed the method in the following.
@@
 Static evaluation on the complete logging task was conducted.
-The values in the ‘Original’ lines indicate the reported results from the respective methods’ own evaluations.
+The values in the Original lines indicate the reported results from the respective methods’ own evaluations.
 The best performance evaluated by AL-Bench across different metrics is highlighted.
@@
 Although $UniLog_{d}s$ is equipped with a more powerful backbone, it still struggles to choose the appropriate position for the log statement.
-As widely demonstrated that the LLMs are not good at counting numbers (Ahn et al., 2024), we might change the output format by directly adding log statement into the source code rather than generate the exactly line number.
-Furthermore, by simply analyzing the control flow graph of the code, we might be able to exclude positions where logging is not feasible to leave less choices.
+As widely demonstrated that the LLMs are not good at counting numbers (Ahn et al., 2024), we might change the output format by directly adding log statements into the source code rather than generating the exact line number.
+Furthermore, by simply analyzing the control flow graph of the code, we might be able to exclude positions where logging is not feasible to leave fewer choices.
 Answer to RQ1.
@@
 Additionally, analyzing the control graph to exclude impossible positions might be useful to enhance log position determination.
-Table 4. Compilation Failure Rates across methods
+Table 4. Compilation Failure Rates Across Methods.
 Methods Compilation Success Rates
@@
 $LEONID_{S}$ 16.4%
-Table 5. Compilation Failure Reasons Analysis
+Table 5. Compilation Failure Reasons Analysis.
 Failed Reason Failed Number
@@
 4.3. RQ2: How well can the predicted log statements be compiled successfully?
-To evaluate tools’ ability to generate compilable log statements, we replace existing logs with predicted ones in the dynamic evaluation dataset and recompile the project to check for successful compilation.
+To evaluate the tools’ ability to generate compilable log statements, we replace existing logs with predicted ones in the dynamic evaluation dataset and recompile the project to check for successful compilation.
 As shown in Table 5, the best-performing tool, FastLog, achieves a 79.9% Compilation Success Rate, followed by $UniLog_{c}l$ at 70.3%, $UniLog_{d}s$ at 60.2%, LANCE at 49.4%, $LEONID_{M}$ and $LEONID_{S}$ at 25.0% and 16.4%, respectively.
@@
 ultimately limiting tool performance in practical applications.
-Table 6. The Semantic Similarity of Logs Printed by the Predicted Log Statements Across Methods.
+Table 6. The semantic similarity between logs printed by the predicted log statements and the ground truth.
 Method Cosine Similarity BLEU BLEU-1 BLEU-4 ROUGE-1 ROUGE-L
@@
 Many predictions record redundant information, while others miss key details.
-The missing key information is primarily due to the prediction being placed beyond the execution path during important events, while setting higher verbosity level in the log statements leads to redundancy.
+The missing key information is primarily due to the prediction being placed beyond the execution path during important events, while setting a higher verbosity level in the log statements leads to redundancy.
 This result highlights that automatic logging tools still have significant room for improvement.
@@
 To minimize inconsistencies, we adapted the released models from previous work (Xie et al., 2023; Mastropaolo et al., 2022, 2023) and sought guidance from the authors of the closed-source model UniLog (Xu et al., 2024) to reproduce the tool under supervision.
-  • External Validity: In this work we evaluate the tools with 10 widely recognized Java projects in different domains.
-The results of each tools are examined in different requirements of logging practices.
+  • External Validity: In this work, we evaluate the tools with 10 widely recognized Java projects in different domains.
+The results of each tool are examined in different requirements of logging practices.
 The novel evaluation method we propose is not limited to Java projects, we will explore its effects in other programming languages in the future.

### v3

@@
 Recently, language model-based techniques have been developed to automate log statement generation based on input code.
-These tools show promising results in their own evaluation.
-However, current evaluation practices in log statement generation face significant challenges.
-The lack of a unified, large-scale dataset forces studies to rely on ad-hoc data, hindering consistency and reproducibility.
-Additionally, assessments based solely on metrics like code similarity fail to reflect real-world effectiveness.
-These limitations underscore the need for a comprehensive public benchmark to standardize evaluation.
-This paper introduces AL-Bench, a comprehensive benchmark designed specifically for automatic logging tools.
-AL-Bench includes a high-quality, diverse dataset collected from 10 widely recognized projects with varying logging requirements and introduces a novel dynamic evaluation approach.
-Different from the existing evaluations that focus only on components of log statements like code similarity, AL-Bench assesses both the compilability of the code with inserted log statements and the effectiveness of the logs generated by them during runtime, which we believe can better reflect the effectiveness of logging techniques in practice.
-AL-Bench reveals significant limitations in the state-of-the-art tools.
-The codes with log statements generated by the state-of-the-art tools fail to compile in 20.1%-83.6% cases.
-In addition, even the best-performing tool only achieves 0.213 cosine similarity between the runtime logs produced by the generated log statements and the ground-truth log statements.
-The results reveal substantial opportunities to further enhance the development of automatic logging tools.
-Logging, Software Maintenance, Logging Benchmark
+While these tools show promising results in prior studies, the fairness of their results comparisons is not guaranteed due to the use of ad hoc datasets.
+In addition, existing evaluation approaches exclusively dependent on code similarity metrics fail to capture the impact of code diff on runtime logging behavior, as minor code modifications can induce program uncompilable and substantial discrepancies in log output semantics.
+To enhance the consistency and reproducibility of logging evaluation, we introduce AL-Bench, a comprehensive benchmark designed specifically for automatic logging tools.
+AL-Bench includes a large-scale, high-quality, diverse dataset collected from 10 widely recognized projects with varying logging requirements.
+Moreover, it introduces a novel dynamic evaluation methodology to provide a run-time perspective of logging quality in addition to the traditional static evaluation at source code level.
+Specifically, AL-Bench not only evaluates the similarity between the oracle and predicted log statements in source code, but also evaluates the difference between the log files printed by both log statements during runtime.
+AL-Bench reveals significant limitations in existing static evaluation, as all logging tools show average accuracy drops of 37.49%, 23.43%, and 15.80% in predicting log position, level, and message compared to their reported results.
+Furthermore, with dynamic evaluation, AL-Bench reveals that 20.1%-83.6% of these generated log statements are unable to compile.
+Moreover, the best-performing tool achieves only 21.32% cosine similarity between the log files of the oracle and generated log statements.
+These results underscore substantial opportunities to advance the development of automatic logging tools.
+We believe this work establishes a foundational step in furthering this research direction.
+Software Maintenance, Logging, Benchmark
 copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Software and its engineering
@@
 As software grows in size and complexity, logging has become increasingly essential to ensuring software reliability (He et al., 2021; Chen and Jiang, 2021).
-Logging means writing log statements into the source code, which generate runtime logs that record valuable information for a range of downstream tasks such as anomaly detection (Zhang et al., 2019; Lou et al., 2010; He et al., 2016; Du et al., 2017; Zhang et al., 2019), fault diagnosis (Zou et al., 2016), root cause analysis (Amar and Rigby, 2019; He et al., 2018b; Lin et al., 2016), and program verification (Ding et al., 2015; Shang et al., 2013).
+Logging means writing log statements into the source code, which generate runtime logs that record valuable information for a range of downstream tasks such as anomaly detection (Zhang et al., 2019; Lou et al., 2010; He et al., 2016; Du et al., 2017; Zhang et al., 2019), fault diagnosis (Zou et al., 2016; Xu et al., 2023, 2024c), root cause analysis (Amar and Rigby, 2019; He et al., 2018b; Lin et al., 2016; Xu et al., 2025), and program analysis (Ding et al., 2015; Shang et al., 2013; Xu et al., 2024b).
 The effectiveness of these downstream tasks heavily relies on the quality of the software logs (He et al., 2018a).
@@
 Figure 1. An example of logging statement generation. Logging statement generation can be separated as three parts: determining the position, selecting the level, and specifying the message.
-Numerous logging tools have been developed to assist software developers by automatically suggesting log statements based on provided code snippets (Mastropaolo et al., 2022; Liu et al., 2021; Xu et al., 2024; Xie et al., 2023).
+To help software developers implement effective and efficient logging practices, researchers have been focusing on exploring methodologies for automatic logging in recent years.
 As illustrated in Figure 1, automatic logging typically includes three steps: (1) determining the position, (2) deciding the verbosity level, and (3) specifying the message to be recorded.
-Leveraging the advanced text generation capabilities of large language models (LLMs) (Floridi and Chiriatti, 2020; Raffel et al., 2020), LANCE (Mastropaolo et al., 2022) is the first end-to-end tool to integrate positioning, level selection, and message generation.
-UniLog (Xu et al., 2024) employed a warm-up and in-context learning (ICL) strategy to enhance performance.
-FastLog (Xie et al., 2023) improved the generation efficiency while maintaining precision.
+To address the automatic logging tasks, numerous logging methods have been proposed thanks to the advanced development of the Large Language Model (LLM).
+For example, LANCE (Mastropaolo et al., 2022) was introduced as the first end-to-end logging method, seamlessly integrating the recommendation of log statement position, verbosity level, and message content.
+Building on this foundation, UniLog (Xu et al., 2024a) employed a warm-up and in-context learning (ICL) strategy to enhance performance.
+FastLog (Xie et al., 2024) improved the generation efficiency while maintaining precision.
 LEONID (Mastropaolo et al., 2023), based on LANCE, combined with deep learning and information retrieval technologies to enhance performance.
 SCLogger (Li et al., 2024) adapted static analysis to extend the context for the code snippet.
-These studies typically evaluate performance using ad-hoc data splitting from the entire dataset, focusing on metrics such as the accuracy of log statement components, including position, level, and message.
-Machine translation metrics, such as BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004), are also employed to evaluate the quality of generated log messages.
-Although these evaluations offer valuable insights into logging tool performance, the use of low-quality data and incomplete assessments undermines the reliability of the results.
-First, loose standards collection and inappropriate clean strategy compromise the quality of evaluation data and effectiveness in assessing the performance.
-Previous evaluation datasets were typically created by splitting the entire dataset.
-The data selection rules of the entire dataset are commonly loose standards to ensure sufficient data for training.
-These criteria fail to ensure the quality and consistency of the data.
-Moreover, to accommodate the limitations of the tools, they filtered out all instances exceeding 512 tokens, ignoring the long code snippets that are commonly encountered in real-world development environments.
-This approach undermines the overall effectiveness and real-world applicability of the evaluation results, as it does not accurately reflect the true complexity of software projects.
-Second, the current evaluation method does not verify whether the generated log statements are compilable.
-Generating the compilable log statements is the first requirement when applying automatic logging tools in practice.
-To relieve developers from the heavy effort required to design and maintain log statements (Chen and Jiang, 2019, 2017), the basic requirement is to ensure that our tool can be seamlessly integrated into the DevOps process without introducing additional errors that require extra debugging effort from developers.
-Current evaluation methods merely focus on whether each component of log statements (i.e., position, verbosity level, message) matches the ground truth but cannot assess whether the generated log statements might introduce compilation errors such as wrong code format or use undefined variables.
-Evaluating the compilability of predicted log statements reflects the effectiveness and reliability of logging tools, which are essential for practical use.
-Third, the evaluation method cannot evaluate the quality of runtime logs generated by the predicted log statements.
-Current evaluation methods assess the performance of tools based on the correctness of individual components of log statements.
-However, current metrics struggle to accurately reflect the quality of runtime logs in the real execution environment.
-Even a slight shift of log statement can lead to the miss of essential runtime details, such as a several-line shift from the ground truth or a mismatch in verbosity level.
-For instance, a minor difference in verbosity levels (e.g., debug vs.
-info) can cause critical information missed due to log level threshold settings in the source code.
-Therefore, we need to evaluate the quality of log statements in a real execution environment, with the goal of obtaining appropriate logs in specific scenarios, rather than merely focusing on the correctness of individual components or relying on statistical metrics to reflect the performance of logging tools.
-Our Work.
-To address these challenges, we introduce AL-Bench: a comprehensive benchmark featuring a large-scale, diverse dataset and a novel approach for evaluating both the compilability and runtime logs produced by generated log statements.
-Our dataset comprises 42,224 instances collected from 10 popular, high-quality GitHub projects (GitHub, [n.
-d.]) spanning different domains with varying logging requirements, providing a robust foundation for evaluating automatic logging methods.
-Beyond the dataset, AL-Bench introduces a novel evaluation method called dynamic evaluation, which involves reintegrating generated log statements into real project code, followed by recompiling and executing them.
-This process allows for a realistic assessment of both compilability and runtime logs, highlighting major limitations in state-of-the-art tools: even the best tools fail to compile in 20.1% of cases, and the logs produced by generated log statements show only 0.213 cosine similarity to ground-truth logs.
-Through its rigorous and standardized evaluation approach, AL-Bench bridges the gap between real-world logging requirements and prior assessments, highlighting substantial opportunities to further advance the development of automatic logging tools.
-This paper’s contributions are summarized as follows:
+However, current logging research faces a significant challenge: The absence of a standardized evaluation dataset and methodology hinders fair comparisons between different logging methods.
+This issue manifests in two key aspects: (1) The test sets of different logging methods are often derived from their own training data, leading to inconsistent experimental performance and hindering fair comparisons.
+(2) The evaluation metrics used for different logging methods are not fully consistent, resulting in varying advantages across different dimensions.
+Given these challenges, we argue that establishing a comprehensive logging benchmark is essential.
+However, it is infeasible to naively construct a benchmark via ensembling existing evaluation data and metrics due to two major limitations of existing evaluation datasets and methods:
+First, existing evaluation datasets lack the quality required for reliable assessment.
+For example, 18.45% cases (2218/12020) in LANCE’s test set and 30.36% cases (2197/7237) in FastLog’s test set involve numerous bad patterns of log messages.
+As illustrated in Figure 2 these anomalies predominantly fall into the following types: (a) duplicated content, (b) empty string, (c) duplicated empty and special characters, (d) contents that mismatch the log levels, and (e) explicit type casting.
+These issues distort evaluation outcomes, as tools producing such flawed log statements tend to appear better on metrics despite having poor logging quality.
+Moreover, these datasets are restricted to cases with very short contents ($\leq$ 512 tokens), a limitation imposed by their use of foundation models (e.g., T5, PLBART), which support a maximum input of 512 tokens.
+This constraint not only reflects a compromise to accommodate their models but also inhibits the evaluation of realistic, function-level logging capabilities, as many industrial functions exceed this token limit.
+Consequently, the datasets fail to meet the requirements for assessing logging quality in real-world scenarios.
+Second, current evaluation methods lack a fine-grained view for comprehensive assessment.
+Existing evaluations primarily focus on static code similarity between candidate log statements generated by logging methods and oracle log statements written by developers.
+However, the code compilability is entirely neglected and can not be measured by static evaluation.
+Our preliminary analysis revealed that 20%-80% of log statements generated by current methods fail to compile, a figure that has not been previously reported.
+Furthermore, such static similarity fails to capture logging performance during program execution.
+For instance, placing a log statement inside or outside a loop may differ by only one line in code but result in vastly different numbers of log entries at runtime.
+Conversely, moving a log statement above or below a comment may have no effect on the runtime logs despite a one-line difference in the code.
+Such runtime differences are overlooked in static evaluations.
+These limitations highlight fundamental flaws in the current evaluation approach.
+Figure 2.
+Bad patterns in existing datasets (using instances from LANCE dataset as examples): (1) Duplicated Variable records the same information multiple times, costing redundant overhead for both printing and recording logs in runtime.
+(2) Empty String provides insufficient context, hindering the effectiveness of debugging through printed logs.
+(3) Unpredictable Character contains numerous meaningless special tokens, making both the logging methods hard to predict and their printed logs difficult to parse for downstream analysis (Adzic, 2006).
+(4) Wrong Verbosity Level often misleads developers for debugging and fault localization (Chen and Jiang, 2017).
+(5) Explicit Cast couple logs to variable type casting, might cause runtime type conversion errors and system crash (Chen and Jiang, 2017).
+To address these two challenges, our key insights are: (1) Log statements from the most widely used and frequently updated repositories tend to be of higher quality with fewer bad logging patterns.
+(2) Log files printed in runtime can serve as a proxy for evaluating log statements quality with a more comprehensive view of program logging performance.
+Based on these insights, we introduce AL-Bench, the first unified logging benchmark involving a large-scale evaluation dataset and a comprehensive evaluation method for assessing both the static log statements and runtime log files printed via code execution.
+Our dataset comprises 21,804 instances mined from the 10 most popular and actively maintained GitHub projects (GitHub, [n.
+d.]), each with at least 10,000 stars, and 500 log-related issues spanning diverse domains and application scenarios.
+Furthermore, AL-Bench introduces dynamic evaluation in addition to traditional static evaluation for log statements.
+Specifically, it not only applies a suite of static metrics to quantify the quality of generated log statements but also reintegrates them into real project code, compiles, and executes them.
+This approach enables a realistic assessment beyond the traditional static approach, highlighting major limitations in existing logging methods: even the best-performing logging method fails to compile in 20.1% of cases, and the logs printed in runtime show only 21.32% cosine similarity to the oracle logs.
+By providing a unified, standardized, and comprehensive evaluation dataset and methodology, we believe AL-Bench establishes a foundational step toward advancing research in automatic logging.
+To sum up, our contributions are shown as follows:
   1. (1)
-    We collected a high-quality, diverse, and large-scale dataset comprising 42,224 instances from 10 popular, high-quality GitHub (GitHub, [n.
+    We collected a high-quality, diverse, and large-scale dataset comprising 21,804 instances from 10 popular, high-quality GitHub (GitHub, [n.
 d.]) projects, spanning various domains with differing logging requirements.
   2. (2)
-    We propose a novel dynamic evaluation approach that assesses generated log statements in real-world settings by reintegrating them into projects to evaluate their compilability and resulting runtime logs.
+    We propose a dynamic evaluation approach that reintegrates the generated log statements into projects to evaluate their printed runtime logs as a more fine-grained proxy for illustrating logging performance.
   3. (3)
@@
     All the data and code for AL-Bench are publicly available [^1] https://github.com/shuaijiumei/logging-benchmark-scripts, providing valuable resources for both developers and researchers to advance the field of automatic logging.
-2. background and motivation
-This section introduces the overview of the current research on automatic logging tools, followed by the shortcomings in the evaluation work, and explains the motivation for AL-Bench.
-2.1. Log Statement Generation
+2. Background
+2.1. Related Work of Automatic Logging
 Logging, the process of generating informative log messages with appropriate verbosity levels at strategically placed locations within code, has long been recognized as a critical challenge in software engineering (He et al., 2021; Chen and Jiang, 2019, 2021; Chen et al., 2021).
@@
 However, none of them can generate a complete log statement.
-Recently, Mastropaolo et al.
-(Mastropaolo et al., 2022) proposed the first end-to-end tool LANCE to generate complete log statements based on T5 (Raffel et al., 2019).
-Following LANCE, Xu et al.
-(Xu et al., 2024) proposed UniLog to adapt ICL and warm-up strategy to enhance the LLM ability for generating log statements.
-Xie et al.
-(Xie et al., 2023) proposed FastLog increase the generation time while keeping the accuracy, and Mastropaolo et al.
-(Mastropaolo et al., 2023) further proposed LEONID with a combination of Deep Learning (DL) and Information Retrieval (IR) achieving a better performance.
-While end-to-end automatic logging tools have demonstrated promising results in their respective evaluations, our analysis reveals notable issues in both the datasets used and the evaluation methods employed.
-Figure 2. The low-quality examples in the LANCE dataset follow the strictest data collection rules.
-2.2. Limitations of Evaluation Methodology
-2.2.1. Evaluation Data
-Current evaluation datasets are ad-hoc, derived from splitting the entire collected data according to loose standards.
-More specifically, to ensure a sufficient amount of training data, previous studies had to adopt lenient rules for data collection.
-Additionally, some tools, to accommodate the limitations of their backbone models, further filter out data exceeding 512 tokens in length.
-These strategies compromise data quality, reducing the reliability and effectiveness of evaluation results.
-For example, LANCE adopted the strictest data collection rules among all tools, requiring a minimum of 500 commits, 10 contributors, and 10 stars, while also excluding forked repositories on GitHub (GitHub, [n.
+Recently, with the advanced capabilities of large language models (LLMs) (Floridi and Chiriatti, 2020; Raffel et al., 2020), numerous LLM-based logging methods have been proposed.
+Specifically, LANCE (Mastropaolo et al., 2022) was introduced as the first end-to-end logging method, seamlessly integrating the recommendation of log statement position, verbosity level, and message content.
+Building on this foundation, UniLog (Xu et al., 2024a) employed a warm-up and in-context learning (ICL) strategy to enhance performance.
+FastLog (Xie et al., 2024) improved the generation efficiency while maintaining precision.
+LEONID (Mastropaolo et al., 2023), based on LANCE, combined with deep learning and information retrieval technologies to enhance performance.
+SCLogger (Li et al., 2024) adapted static analysis to extend the context for the code snippet.
+2.2. Philosophy of Logging Evaluation
+Logging quality evaluation has always been a critical yet challenging problem in automatic logging (Bogatinovski et al., 2022; Li et al., 2019; Chen and Jiang, 2017, 2021; He et al., 2021).
+In an ideal scenario, high-quality logs should accurately capture key runtime state information of a software system while avoiding excessive logging that may lead to unnecessary consumption of computational and storage resources.
+Intuitively, the most direct approach to assessing logging quality is to evaluate how well the logs support specific downstream tasks (Yuan et al., [n.
 d.]).
-However, even in the LANCE dataset, low-quality data is pervasive.
-As shown in Figure 2, several patterns of low-quality logging practices persist.
-These patterns include the duplication of variables within a single log statement, creating unnecessary redundancy; the inclusion of empty strings or meaningless content, resulting in uninformative messages; excessive use of special characters or punctuation, making printed logs difficult to parse; and mismatched logging levels, where critical messages are assigned inappropriate severities, leading to misclassification (Adzic, 2006; Chen and Jiang, 2017).
-Using low-quality data as an evaluation dataset fails to produce reliable results.
-Such results do not align with the expectation that logging tools should generate high-quality log statements within the code.
-Furthermore, to accommodate the input length limitations of these tools, previous studies (Mastropaolo et al., 2022, 2023; Xu et al., 2024; Xie et al., 2023) have filtered out instances longer than 512 tokens.
-While this data-cleaning strategy reduces the complexity of evaluation data, it also introduces significant biases by excluding longer code snippets, which are common in real-world development environments.
-This approach simplifies the evaluation process but fails to capture the full scope of challenges that logging tools would face when dealing with complex and extended code bases.
-Consequently, the evaluation results may not accurately reflect the tools’ ability to handle the demands of real-world software development, where longer snippets and more intricate code structures are prevalent.
-Therefore, there is a pressing need for a public, large-scale, high-quality, and diverse benchmark dataset that can better represent real-world codebases and provide a standardized platform for evaluating automatic logging tools.
-Figure 3.
-An example of the limitation of static evaluation, where only the quality of log statements is considered.
-It cannot assess the compilability of the code or quality of the printed runtime logs.
-2.2.2. Evaluation Method
-The current evaluation method focuses on assessing model performance by comparing the accuracy of each log statement component (i.e., position, verbosity level, message).
-We refer to this as static evaluation in this paper.
-Figure 3 provides an example of static evaluation to demonstrate how it works.
-In this example, only the verbosity level matches, so the matched level count will increase by one, and BLEU scores will be used to calculate the average score.
-Finally, the matched portion of each component and the average BLEU score will be combined to reflect the performance of the tools.
-These statistical metrics provide insights into evaluating the quality of predicted log statements; however, this approach still does not align with the goal of ensuring that the predicted log statements generate high-quality logs.
-Figure 4. An example of compilation failure: Unreachable Statements
-First, it cannot assess the compilability of predicted log statements.
-Compilability is a fundamental requirement for generating appropriate logs.
-This capability ensures that log statements can be applied to the source code without introducing compilation errors, avoiding wasting developers’ efforts to debug for the predicted log statements.
-However, the current evaluation methodology, which focuses solely on comparing the individual components of log statements, cannot identify predicted log statements that use undefined variables or contain incorrect code syntax which could introduce compilation errors.
-As shown in Figure 4, the predicted log statement is injected into an unreachable position.
-Applying static evaluation would fail to recognize that this prediction could cause compilation errors.
-Instead of classifying it as a completely incorrect prediction, the evaluation would count it as a level match instance and assign it a BLEU score of 55.03, contributing to the statistical value.
-This could introduce significant bias when evaluating the performance, appearing favorable in metrics but performing poorly in actual usage.
-Second, it cannot assess the runtime logs generated by predicted log statements.
-The goal of writing log statements is to obtain the appropriate logs during software execution so that the quality of log statement is ultimately decided by the quality of runtime logs.
-Although comparing the correctness of log statement components provides insights into assessing the quality of logs, it can introduce bias when using statistical metrics that focus solely on log statements to evaluate the performance of logging tools.
-A simple example is a mismatched verbosity level between the predicted log statement and the ground truth.
-In a case where only the verbosity level is mismatched while the other components are exactly matched, this instance would receive a high score in static evaluation.
-However, in a real-world scenario, for important error logs, if the predicted level is below the threshold, the logs will not be generated, regardless of how critical the message is.
-Conversely, for debug logs, if the predicted level exceeds the threshold, these logs could be printed in the production environment, leading to issues such as sensitive information leakage and increased storage costs.
-These two situations demonstrate that even a shift in verbosity level can turn a log statement into a completely bad case.
-A more complicated case is presented in Figure 5.
-Under static evaluation, this prediction with only a shift in two lines in position, also could be rewarded highly in statistic metrics.
-However, it is completely incorrect: placing the log statement at line 2 risks a NullPointerException if the timer object is null when timer.get() is called.
-This would cause the program to crash during execution, meaning that not only would the logs fail to be generated, but the entire program could terminate unexpectedly.
-In real-world scenarios, this kind of error is critical and goes unnoticed in static evaluations.
-While static evaluation rewards this prediction for structural correctness, it fails to account for the actual behavior of the code when executed.
-Therefore, this example reinforces the need for more comprehensive evaluation methods that go beyond code-level metrics and consider the real-world implications of log statements.
-In conclusion, a log statement’s quality should be determined by its ability to avoid introducing extra errors and generate meaningful, contextually appropriate logs.
-Figure 5.
-An example of the limitation of static evaluation.
-The prediction differs from the ground truth by a slight upward shift of two lines.
-It causes a NullPointerException when timer is null and fails to generate logs.
-Insights: The motivating study underscores major limitations in current evaluation efforts, including the lack of a diverse, high-quality dataset and methods that do not meet real-world needs.
-To improve log statement evaluation, we should collect a high-quality, diverse dataset and use an execution-based method to directly evaluate the runtime logs generated by log statements.
+However, using downstream tasks as proxies to measure general-purpose logging quality in a scalable way is both theoretically and practically problematic for several reasons:
+  1. (1)
+    First, it is very challenging to cover all potential downstream tasks for logs exhaustively.
+There are countless log-based downstream tasks: beyond the tasks with clear task formulations (e.g., debugging, anomaly detection, root cause analysis), logs are also used for a wide range of tasks that lack a universal formulation (e.g., recording sequences of runtime events to monitor system status), often driven by the unique needs of each system’s developer (Zhang et al., 2019; Xu et al., 2024c; Amar and Rigby, 2019).
+Even if we only choose a fixed set of tasks with clear formulations to evaluate logging quality, these proxies can only assess task-specific aspects rather than provide a comprehensive measure of overall logging quality.
+This limitation makes them unreliable indicators of general-purpose logging quality.
+  2. (2)
+    Second, it is not practical to build testbeds for these downstream tasks across all software systems.
+While it might be theoretically possible to create a testbed for evaluating logging quality in common tasks like anomaly detection, assessing tasks like whether logs capture key information for event monitoring needs is challenging, as it depends heavily on subjective developer requirements.
+Moreover, even creating fault datasets to support log-based anomaly detection is also challenging since different software systems have unique fault characteristics, making scalable fault injection across diverse environments infeasible (Zhao et al., 2021).
+  3. (3)
+    Third, the trade-off between logging effectiveness and efficiency is another unmeasurable aspect of logging quality.
+Massive logging offers detailed runtime data for fault diagnosis but increases storage costs and resource usage, while sparse logging reduces expenses but risks missing critical information.
+System designers balance these tradeoffs primarily through operational experience from maintaining and debugging specific systems.
+However, such experiences are inherently context-dependent, varying across systems due to divergent functional, architectural, and operational priorities (Rong et al., 2023).
+Consequently, it remains challenging to establish a universal trade-off mechanism for assessing the appropriateness of logging strategies for each specific software.
+As an alternative, logging research often adopts a similarity-based approach to evaluate logging quality.
+The core idea is that, since assessing log quality via downstream tasks is challenging, we can use high-quality log statements as references to evaluate new ones.
+Thus, by constructing benchmarks, we can check if a logging method generates log statements similar to these reference ones, making evaluation scalable with log statements from open-source software.
+However, this similarity-based approach does have limitations.
+It may not accurately assess log statements of higher quality than the reference, as they could receive lower similarity scores simply for not matching closely enough.
+To mitigate this, reference log statements must be of sufficiently high quality, which poses a challenge in data collection.
+We argue that log statements from the most widely-used, time-tested projects are the best reference candidates.
+Additionally, current evaluations focus solely on static code, but small changes in code can lead to significant differences in log content.
+Therefore, evaluating both static code and the generated log files provides a more comprehensive measure of logging quality.
+These insights inform the development of our first logging benchmark: AL-Bench.
 3. AL-Bench
-In this section, we introduce our benchmark, AL-Bench, which builds on insights from previous studies (Mastropaolo et al., 2022, 2023; Xu et al., 2024; Xie et al., 2023).
-AL-Bench consists of two key parts: static evaluation and dynamic evaluation, both of which we will introduce in detail.
-The static evaluation component includes a high-quality dataset with 42,224 instances and five metrics, based on the static evaluation method outlined in Section 2.2.2.
-The dynamic evaluation component introduces a separate dataset with 2,238 instances and four metrics, following the dynamic evaluation approach, which will be described later in this section.
+AL-Bench is a novel benchmark designed to evaluate automatic logging tools in codebases systematically.
+The benchmark comprises 21,804 code snippets and 39,600 log statements extracted from 10 most popular open-source repositories, selected for their active maintenance, diverse application domains, and representation of modern software engineering practices.
+AL-Bench employs a dual evaluation framework: Static Evaluation evaluates the similarity of log statements (e.g., position, level, variable, message) between predicted and reference log statements, while Dynamic Evaluation assesses code compileability and the similarity of logs printed by those log statements during runtime execution.
+Both evaluation methods will be described later in this section.
+3.1. Dataset Construction
+As discussed in Sec.
+1, the quality of the evaluation dataset is crucial for assessing the performance of tools.
+Building on empirical evidence that GitHub repository stars correlate with code quality (Jiang et al., 2024), we determine dataset inclusion criteria for log statement analysis by requiring repositories to have $\geq$ 10,000 stars and $\geq$ 500 log-related issues to ensure community validation and logging relevance.
+Candidate repositories are then ranked by total log statement count to prioritize those with extensive logging practices.
+From the top of this ranked list, we manually curate a final dataset of 10 repositories to ensure that the set covers different priorities in the logging effectiveness-efficiency trade-off, including performance-critical systems (requiring minimal logs) and trace-intensive applications (requiring detailed logs).
+After that, we ensure that all of these repositories are actively maintained, with frequent commits (¿50/month), a high issue resolution rate (¿70%), and recent updates (within the past 6 months).
+This dual-phase approach balances the quantitative scale with domain representativeness.
+As shown in Table 1, our final dataset includes projects with a total of 21,804 code snippets and 39,600 log statements, covering a wide range of logging needs and practices.
+The dataset spans multiple domains, including database management, task scheduling, distributed storage, messaging systems, and IoT platforms.
+These repositories have distinct logging objectives (targeted downstream tasks) based on the specific needs of their respective scenarios and exhibit unique effectiveness-efficiency trade-offs (more logging vs less logging), as shown in Sec.
+2.2.
+For example, database systems like DBeaver minimize logging overhead to sustain performance, while task schedulers such as DolphinScheduler emphasize dependency-tracing logs for runtime monitoring.
+Distributed systems (e.g., Hadoop) enforce rigorous logging for fault tolerance and distributed coordination, contrasting with messaging platforms like Kafka, which track message flows to ensure delivery reliability.
+IoT solutions (e.g., ThingsBoard) leverage logs for real-time device connectivity oversight, whereas security-focused systems like Keycloak prioritize log anonymization to safeguard sensitive data.
+These diverse logging requirements not only demonstrate the heterogeneity and inherent contradictions in logging practices (e.g., performance-sensitive scenarios demand minimal logging overhead, while debugging scenarios necessitate comprehensive event records) but also substantiate the impossibility of establishing universal standards for assessing log code quality.
+By documenting real-world logging practices from community-vetted, widely-used projects across domains, we establish a foundation for evaluating logging strategies in their native environments.
+This domain-rooted dataset supports comprehensive analysis of automated logging techniques across diverse operational scenarios.
+Table 1. Details of AL-Bench dataset.
+Dataset Domain # Code w/ LogStmt # LogStmt
+Dbeaver Database Management 1,178 1,707
+Dolphinscheduler Task Scheduling 883 1,855
+Doris High Performance Database 1,635 2,860
+Flink Data Processing 1,916 3,150
+Hadoop Distributed Storage 8,351 14,987
+Kafka Messaging Systems 1,667 3,308
+Keycloak Identity and Access Management 491 926
+Pulsar Messaging Systems 3,279 6,347
+Thingsboard IoT Platform 1,544 2,588
+Zookeeper Distributed Coordination 860 1,872
+Total - 21,804 39,600
 3.2. Static Evaluation
-3.1.1. Static Evaluation Method.
-Static evaluation focuses on log statement components—position, verbosity, and message—and has been the primary method in prior studies (Mastropaolo et al., 2022, 2023; Xu et al., 2024; Xie et al., 2023), with details provided in Section 2.2.2.
-3.1.2. Dataset Construction
-As discussed in Section 2.2.1, the quality of the evaluation dataset is crucial for assessing the performance of tools.
-The strictest existing rules for dataset collection—500 commits, 10 contributors, and 10 stars—are inadequate to ensure the quality of the evaluation dataset.
-Drawing inspiration from the use of GitHub repository stars as an effective metric for identifying high-quality code datasets  (Jiang et al., 2024), and considering the unique characteristics of log statements, we propose three critical criteria for dataset selection: repositories must have at least 10,000 stars, 1,000 log statements, and 500 log-related issues to qualify for inclusion in our dataset.
-These standards guarantee well-maintained, widely used projects where developers prioritize high-quality log statements.
-Given the diverse requirements for log statements in different scenarios, we ultimately selected 10 projects based on domain diversity, industrial applications, and their relevance to prior logging studies (Chen and Jiang, 2017, 2019; Li et al., 2018, 2019; Ding et al., 2022, 2023; Li et al., 2020).
-As shown in Table 1, our final dataset includes projects with a total of 22,787 code snippets and 42,224 log statements, covering a wide range of logging needs and practices.
-The dataset spans multiple domains, including database management, task scheduling, distributed storage, messaging systems, and IoT platforms, ensuring diversity in logging scenarios.
-Each domain presents unique requirements for log statements.
-For example, database management systems such as DBeaver and Doris prioritize minimizing the impact of logging on high performance.
-Task scheduling systems, including DolphinScheduler, rely on logging to trace task dependencies and monitor runtime statistics.
-Similarly, distributed systems like Hadoop and Zookeeper require robust logging practices to address challenges in distributed coordination, fault tolerance, and scalability, which differ from the requirements of other command logs.
-Messaging systems such as Kafka and Pulsar have adapted logging practices to trace message flows, ensure reliable message delivery, and debug asynchronous communication.
-Meanwhile, IoT platforms like ThingsBoard utilize logging to manage device connectivity, monitor data streams, and enable real-time system oversight.
-Identity and access management systems such as Keycloak prioritize protecting sensitive information in logs to prevent privacy breaches.
-By including diverse projects across these domains, the dataset ensures comprehensive coverage of different logging practices, making it suitable for benchmarking and analyzing log-related tasks under various operational contexts.
-In addition to emphasizing data quality, we also addressed the potential risk of data contamination.
-Since all our data were extracted from public GitHub repositories, which may have been used for training pre-trained models, we implemented precautions to minimize this risk.
-Specifically, we collected the latest version of each project to reduce the likelihood of it being included in any model’s pre-training data.
-Furthermore, we wrapped the code snippets in a common class and standardized the formatting using Google-Java-Format (Google, [n.
-d.]).
-This approach altered the format of the code to prevent pre-trained models from recognizing the same information and structure.
-These strategies have been demonstrated as effective in recent studies (Xu et al., 2024; Tian et al., 2024).
-Although we adopted effective strategies, data contamination cannot be entirely avoided (Cao et al., 2024).
-However, our methods minimize this risk and have been proven effective in previous work, providing a solid foundation for analysis and evaluation.
-In the future, we plan to regularly update our dataset to ensure that the evaluation data remains current.
-After completing the necessary appeal actions, we finalized our static evaluation dataset.
-Table 1. Details of AL-Bench dataset.
-Dataset Domain Code snippets with log statements Total log statements
-Dbeaver Database Management 1,182 1,725
-Dolphinscheduler Task Scheduling 898 1,918
-Doris High Performance Database 1,686 2,965
-Flink Data Processing 1,956 3,268
-Hadoop Distributed Storage 8,751 16,002
-Kafka Messaging Systems 1,706 3,421
-Keycloak Identity and Access Management 533 1,004
-Pulsar Messaging Systems 3,590 7,254
-Thingsboard IoT Platform 1,601 2,731
-Zookeeper Distributed Coordination 884 1,936
-Total - 22,787 42,224
+3.2.1. Task Formulation
+We structure each entry in our dataset as a tuple $<Code_{w/o~{}LogStmt},LogPos,LogStmt>$.
+As illustrated in Figure 3, $Code_{w/o~{}LogStmt}$ represents the input code context with one log statement deliberately removed, $LogPos$ indicates the ground truth position for the missing log statement, and $LogStmt$ represents the exact log statement to be predicted.
+This process yielded 39,600 high-quality instances, providing a robust foundation for evaluating log generation models.
 3.2.2. Metrics
-Building on previous studies (Mastropaolo et al., 2022, 2023; Xu et al., 2024; Xie et al., 2023), We adopted five metrics for static evaluation.
-In addition to the previously established metrics, we propose two new ones: Dynamic Variable Accuracy and Static Text BLEU, designed to better reflect the real-world quality requirements of log statements.
-Metric 1: Level Accuracy (LA): Level Accuracy focuses on log level values, which indicate the importance of the message and are a crucial part of log statements.
-Some common levels, like ”info,” refer to the normal information of runtime behavior.
+We introduce six static evaluation metrics to assess the similarity between the generated and reference log statements.
+Metric 1: Position Accuracy (PA): PA focuses on the precise line of log statements within the source code.
+The correct placement of log statements helps accurately trace the execution flow and diagnose issues.
+For Position Accuracy, we rigorously compare the predicted positions of log statements with their actual positions in the source code.
+This metric is calculated by taking the number of correctly positioned log statements $P_{c}$ and dividing it by the total number of log statements $N_{a}$ to obtain the accuracy value: $PA=\frac{P_{c}}{N_{a}}$.
+Metric 2: Level Accuracy (LA): LA evaluates the exact match between predicted and reference log levels, which are essential for prioritizing operational events in DevOps pipelines.
+Common log levels carry distinct semantic implications, like ”info,” referring to the normal information of runtime behavior.
 ”Warning” indicates potential problems that might not immediately cause disruption but could lead to future issues if not resolved.
 ”Error” refers to runtime anomalies or issues that need to be addressed.
-Each level refers to different meanings so for Level Accuracy we strictly compared the level of predictions and source code and used the exactly matched level number $L_{c}$ divided by the total number of log statements $N_{a}$ to get the value of this metric: $LA=\frac{L_{c}}{N_{a}}$.
-Metric 2: Position Accuracy (PA): Position Accuracy focuses on the precise location of log statements within the source code, which is crucial for tracking and debugging software behavior.
-Correct placement of log statements helps in accurately tracing the execution flow and diagnosing issues.
-For Position Accuracy, we rigorously compare the predicted positions of log statements with their actual positions in the source code.
-This metric is calculated by taking the number of correctly positioned log statements $P_{c}$ and dividing it by the total number of log statements $N_{a}$ to obtain the accuracy value: $PA=\frac{P_{c}}{N_{a}}$.
-Metric 3: Message Accuracy (MA): Message Accuracy evaluates how accurately the predicted log messages match the ground truth log messages, which is essential for providing meaningful and relevant information during runtime.
+LA is calculated as the ratio of correctly predicted levels $L_{c}$ to the total log statements $LA=\frac{L_{c}}{N_{a}}$.
+Metric 3: Average Level Distance (ALD): ALD further quantifies the severity deviation of mispredicted levels from the reference level.
+We assign ordinal values to log levels: trace: 0, debug: 1, info: 2, warn: 3, error:4, fatal: 5, and compute the absolute difference between predicted level $L_{p}^{(i)}$ and oracle level $L_{a}^{(i)}$ for each log.
+ALD is the mean deviation across all instances: $ALD=\frac{1}{N_{a}}\sum_{i=1}^{N_{a}}\left|L_{p}^{(i)}-L_{a}^{(i)}\right|$.
+Metric 4: Message Accuracy (MA): MA evaluates how accurately the predicted log messages match the oracle, which is essential for providing meaningful and relevant information during runtime.
 The content of log messages helps developers understand the system’s behavior, and inaccuracies in message generation can lead to confusion or missed insights during debugging.
@@
 This metric is calculated by determining the number of log messages that are fully identical to the ground truth $M_{c}$ and dividing it by the total number of log messages $N_{a}$, yielding the accuracy value: $MA=\frac{M_{c}}{N_{a}}$.
-Metric 4: Dynamic Variable Accuracy (DVA): Dynamic Variable Accuracy focuses on the dynamic variable in log message, for example, (”The server is running, {}”, status), in this message, status is regarded as the dynamic variable in log message which might vary according to the runtime behaviors of software.
-And in the log message (”The server run on the ports, {}”, args.status ?
-localPort : remotePort), the whole expression (args.status ?
-localPort : remotePort) is reckoned as the dynamic variable which could decide the port information recorded in log files.
+Metric 5: Dynamic Expression Accuracy (DEA): DEA evaluates whether generated logs preserve the structural integrity of runtime expressions, including both individual variables and composite logic (e.g., ternary operators, arithmetic).
+For example, in the log template:(”The server run on the ports, {}”, args.status ?
+localPort : remotePort), the conditional expression (args.status ?
+localPort : remotePort) is treated as a single semantic unit.
 We aim to use this metric to ensure that the dynamic information recorded in logs remains consistent.
-We extract the dynamic variables from both the source code and predictions and then compare them to record instances where they match exactly.
-This metric is calculated by taking the number of the exactly matched $DP_{c}$ and dividing it by the total number of log statements $N_{a}$ to obtain the accuracy value: $DVA=\frac{DP_{C}}{N_{a}}$.
-Metric 5: Static Text BLEU (STB): Static Text BLEU focuses on the static part of the log message.
+This metric is calculated by taking the number of the exactly matched $DP_{c}$ and dividing it by the total number of log statements $N_{a}$ to obtain the accuracy value: $DEA=\frac{DP_{C}}{N_{a}}$.
+Metric 6: Static Text Similarity (STS) (with BLEU or ROUGE): STS focuses on the static part of the log message.
 Unlike the dynamic variable, the static part always records the same information in log files, which will not vary due to the runtime behavior of software.
 For example, in the log message (”The server is running, {}”, status), ”The server is running, {}” is regarded as the static part.
-Since this part primarily consists of natural language content, we use BLEU (Papineni et al., 2002) metric to evaluate the quality of the static text.
-In our implementation, we use the BLUE-DM variant (Chen and Cherry, 2014; Shi et al., 2021), i.e., the sentence-level BLEU without any smoothing method.
+Since this part primarily consists of natural language content, we use BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) metrics to evaluate the quality of the static text.
+In our implementation, we use the DM variant of BLEU (Chen and Cherry, 2014; Shi et al., 2021), i.e., the sentence-level BLEU without any smoothing method — coupled with ROUGE-L to holistically assess both lexical precision and long-sequence coherence.
+Specifically, ROUGE-L focuses on the longest common subsequence that effectively captures key operational patterns (e.g., error codes or API call chains) in multi-line logs, while BLEU’s n-gram overlap measurement complements it by evaluating template fidelity at the token level.
+Figure 3.
+$Code_{w/o\ \ LogStmt}$ indicates the code without one log statement, the $LogPos$ means the position of this log statement, $LogStmt$ is the log statement itself.
+Those three structure the evaluation tuple.
 3.3. Dynamic Evaluation
@@
 First, compile the project and run the unit test to obtain ground truth logs.
-Then, replace log statements with predictions, rerun the test to generate new logs, and finally analyze the results.
-3.2.1. Dynamic Evaluation Method
-Different from static evaluation, dynamic evaluation focuses on compiling the code and runtime-generated logs, addressing static evaluation’s inability to verify code compilability and runtime logs.
-To directly assess runtime logs, we generate them using unit tests, which are widely used in software development to verify code functionality in isolated scenarios.
-Unit tests are readily available in most projects and they are designed to test the functionality and behaviors of code when facing different situations.
-They offer a natural method for simulating realistic situations, allowing generating logs without the need for complex runtime environments.
-Figure 6 demonstrates the general workflow of dynamic evaluation.
-The process begins with compiling the source code and executing the unit tests to obtain logs from the original log statements.
-Using tools like Jacoco (Jacoco, [n.
-d.]) and SureFire (surefire, [n.
-d.]), we collect the logs and remove the log statements covered by unit tests in the source code.
-Next, we input the modified source code into an automatic logging tool to generate predicted log statements.
-Then we inserted the predicted log statements back into the source code, replacing the original log statements.
-After recompiling the modified code, we rerun the unit tests to capture the logs produced by the inserted predicted log statements.
-The whole process provides us with two sets of logs—those generated by the original log statements and those generated by the predicted log statements.
-Finally, we evaluate the effectiveness of the predicted log statements using two key metrics: Compilation Success Rate and Log File Similarity.
-Compilation Success Rate ensures that the predicted log statements do not introduce compilation errors, while Log File Similarity measures the similarity between the logs generated by the predicted log statements and those generated by the original log statements.
-In the following sections, we will detail how we built the dataset for dynamic evaluation and introduce the specific metrics used to measure the performance of the automatic logging tools.
-3.2.2. Dataset Construction
-To build the dynamic evaluation dataset, we begin by compiling the entire project to ensure all dependencies are resolved and the project is ready for execution.
+Then, replace log statements with predictions, re-run the test to generate new logs, and finally analyze the results.
+3.3.1. Task Formulation
+In dynamic evaluation, we provide the runtime perspective of evaluating the automatic logging tools.
+Two key evaluation metrics are employed: (1) compilability—ensuring the code with predicted log statements compiles without errors, and (2) log file similarity—measuring the alignment between log files generated by the predicted log statements and the oracle through textual similarity analysis.
+These criteria jointly validate both the functional correctness of log integration and the relevance of logged content.
+As illustrated in Figure 4, the core innovation of our methodology lies in utilizing unit tests to emulate runtime log generation under controlled conditions.
+By executing these tests, we synthetically replicate software runtime states with minimal computational overhead, thereby enabling systematic collection of runtime logs that mirror real-world execution patterns for downstream analysis.
+We define task formulation as a tuple $<Code_{w/o\ \ LogStmt},Code_{w/\ \ LogStmt},Logs>$, where $Code_{w/o\ \ LogStmt}$ is the code snippets without log statements covered by unit test, $Code_{w/\ \ LogStmt}$ is the code snippets with predicted log statement, and $Logs$ is the logs produced by those log statements in $Code_{w/\ \ LogStmt}$ via unit tests.
+If the $Code_{w/\ \ LogStmt}$ is not compiliable, $Logs$ is null.
+To build the dynamic evaluation pair based on our dataset, we begin by compiling the project to ensure all dependencies are resolved and the project is ready for execution.
 Next, we systematically identify all available unit tests within the project.
 For each unit test, we execute it individually while employing the Jacoco Plugin (Jacoco, [n.
-d.]) to trace code coverage, specifically identifying whether the unit test interacts with or covers any log statements in the codebase.
+d.]) to trace code coverage, specifically identifying whether the unit test covers any log statements in the codebase.
 Simultaneously, we use the SureFire Plugin (surefire, [n.
@@
 By correlating Jacoco coverage data with SureFire logs, we can match specific code snippets containing log statements to the corresponding unit tests that cover them, along with the runtime logs they generate.
-This process enables us to construct a comprehensive dataset consisting of triples: the code snippet, the unit test that triggers it, and the recorded logs.
-These triples are critical for dynamic evaluation, as they provide the ground truth for assessing the quality and effectiveness of predicted log statements in actual runtime scenarios.
-To construct a dynamic evaluation dataset, it is essential to select high-quality projects with comprehensive unit tests, as these tests provide a realistic simulation of diverse production environments.
-Each instance of dynamic evaluation requires recompiling the project and executing the test to collect logs, making the process highly time-consuming.
-To balance this intensive time requirement with the need for sufficient dataset diversity and quantity, we employed Hadoop as our dynamic evaluation platform.
-This approach allowed us to build a dataset of 2,238 instances that balances diversity with sufficient size.
-Additionally, we open-source the entire suite of tools used in this evaluation process, enabling researchers and organizations to easily deploy and customize their own dynamic evaluation datasets.
+We finally built 2,238 instances for dynamic evaluation.
 3.3.2. Metrics
-In dynamic evaluation, we proposed four metrics to assess the performance of logging tools: Compilation Success Rate, Log Similarity, False Positive Log Generation Rate, and False Negative Log Generation Rate, which we will introduce below.
-Metric 6: Compilation Success Rate (CSR): Compilation Success Rate measures the syntactic correctness of the predicted log statements.
-As discussed earlier, after replacing the original log statements with the predictions, we recompile the project.
-However, due to issues such as undefined variables in the predictions or missing/outdated dependencies in the project environment, not all predictions can be successfully compiled.
-We recorded the successfully compiled code snippet number as $C_{s}$ and all code snippets as $C_{a}$.
+We propose four dynamic evaluation metrics to assess the similarity of logs printed by the generated and reference log statements in runtime execution.
+Metric 7: Compilation Success Rate (CSR): CSR measures the compilability of the predicted log statements.
+Due to issues such as undefined variables in the predictions or missing/outdated dependencies in the project environment, not all code snippets with predicted log statements can successfully compile.
+We recorded the successfully compiled code snippet number as $C_{s}$ and the evaluation pair number as $C_{a}$.
 The metric is then calculated as: $CSR=\frac{C_{s}}{C_{a}}$.
-Metric 7: Log Similarity Metrics Group (Cosine Similarity, BLEU, ROUGE): This metric evaluates how closely logs generated by predicted statements match those produced by ground truth statements.
-To eliminate unnecessary differences, we remove log headers (e.g., timestamps), retaining only log content.
-For a comprehensive assessment, we apply multiple similarity measures, including Cosine Similarity (Salton et al., 1975), BLEU (Papineni et al., 2002), and ROUGE (Lin, 2004).
+Metric 8: Log File Similarity (LFS) (with COS, BLEU, ROUGE): LFS evaluates how closely logs generated by predicted statements match those produced by ground truth statements.
+We remove log headers (e.g., timestamps) and only assess the main message body of each log entry to eliminate unnecessary differences.
+For a comprehensive assessment, we apply multiple similarity measures, including Cosine Similarity (COS) (Salton et al., 1975), BLEU (Papineni et al., 2002), and ROUGE (Lin, 2004).
 Cosine Similarity, commonly used in text analysis, calculates the cosine of the angle between two TF-IDF (Jones, 2021) vectors, yielding 1 for identical vectors and 0 for orthogonal ones.
@@
 The most commonly used variant is ROUGE-N, which calculates the overlap of n-grams between two texts.
-Metric 8: False Positive Log Generation Rate (FPLR): This metric measures the proportion of predicted log statements that generate logs during unit test execution when the ground truth log statements would not have produced any logs.
+Metric 9: False Positive Log Generation Rate (FPLR): FPLR measures the proportion of predicted log statements that generate logs during unit test execution when the ground truth log statements would not have produced any logs.
 It helps assess whether the predicted log statements introduce unnecessary or redundant logs in scenarios where no log should be generated.
@@
 the metric is calculated as: $FRLR=\frac{FP}{P}$.
-Metric 9: False Negative Log Generation Rate (FNLR): This metric evaluates the proportion of predicted log statements that fail to generate logs during unit test execution when the ground truth log statements should have produced logs.
+Metric 10: False Negative Log Generation Rate (FNLR): FNLR evaluates the proportion of predicted log statements that fail to generate logs during unit test execution when the ground truth log statements should have produced logs.
 It highlights instances where the predicted logs miss important events or information.
@@
 From the perspective of empirical software engineering, we set three research questions:
-  • RQ1: How well are logging tools at predicting log statements in static evaluation?
-  • RQ2: How well can the predicted log statements be compiled successfully?
-  • RQ3: How well do the generated log statements print logs in dynamic evaluation?
-Specifically, RQ1 examines how well logging tools predict log statements using the AL-Bench static evaluation dataset and corresponding static evaluation method.
-This research question aims to evaluate the accuracy and effectiveness of the tools in generating log statements that closely resemble the ground truth.
+  • RQ1: How similar are automatically generated log statements to human-written logging practices?
+  • RQ2: How compilable are the log statements generated by automatic logging tools?
+  • RQ3: How consistent are runtime logs from auto-logged programs with oracle-generated logs?
+Specifically, RQ1 aims to evaluate the performance of the tools in generating log statements that closely resemble the ground truth.
 RQ2 focuses on assessing the compilability of the predicted log statements, determining whether these predictions can be seamlessly integrated into the code without causing compilation errors.
 Finally, RQ3 evaluates the runtime logs produced by the generated log statements.
-We examine not only the generation of inappropriate logs but also the similarity between the predicted logs and the expected logs in cases where logs are correctly generated.
+We examine the similarity between the predicted logs and the expected logs.
 We begin by introducing the automatic logging tools selected for evaluation.
-We will provide a detailed analysis of these tools, focusing on their performance across the three research questions.
-Each tool is assessed for its ability to predict accurate log statements, compile them successfully, and generate meaningful logs.
-This comprehensive evaluation allows us to draw insights into the strengths and limitations of current approaches to automatic logging.
+Then we provide a detailed analysis of these tools by answering the three research questions.
+This comprehensive evaluation allows us to draw insights into the strengths and limitations of current approaches for automatic logging.
 4.1. Automatic Logging Tools
@@
 We reached out to the authors of popular end-to-end automatic logging tools for assistance in rebuilding these tools.
-Because SCLogger (Li et al., 2024) is still under construction, we ultimately selected four methods for evaluation: LANCE (Mastropaolo et al., 2022), LEONID (Mastropaolo et al., 2023), FastLog (Xie et al., 2023), UniLog (Xu et al., 2024).
+Because SCLogger (Li et al., 2024) is still under construction, we ultimately selected four methods for evaluation: LANCE (Mastropaolo et al., 2022), LEONID (Mastropaolo et al., 2023), FastLog (Xie et al., 2024), UniLog (Xu et al., 2024a).
 We detailed the method in the following.
@@
 It takes a method requiring a log statement and outputs a meaningful log message with an appropriate logging level in the correct position.
-Built on the Text-To-Text Transfer Transformer model, $LANCE$ is trained specifically for injecting proper logging statements.
+Built on the Text-To-Text Transfer Transformer (T5) model (Raffel et al., 2019), $LANCE$ is trained specifically for injecting proper logging statements.
 LEONID: $LEONID$ (Mastropaolo et al., 2023) is the updated version of $LANCE$.
@@
 Therefore, we only applied $LEONID_{S}$ for static evaluation.
-UniLog: $UniLog$ (Xu et al., 2024) is the first attempt to adapt Warm-up and In-context-learning strategy to enhance the model’s ability to generate log statements.
+UniLog: $UniLog$ (Xu et al., 2024a) is the first attempt to adapt Warm-up and In-context-learning strategy to enhance the model’s ability to generate log statements.
 Due to limitations in assessing the original UniLog, we reproduced it using two backbone models: CodeLlama-7B (Rozière et al., 2023) and DeepSeek-V3 (DeepSeek-AI et al., 2024).
-We applied the warmup process exclusively to the CodeLlama backbone model, while employing the ICL strategy to construct prompts for both models.
+We applied the warmup process exclusively to the CodeLlama-7B backbone model while employing the ICL strategy to construct prompts for both models.
 The data are sourced from LANCE (Mastropaolo et al., 2022) to warm up and generate ICL content.
@@
 Since LANCE’s data distribution differs from our evaluation data, this may affect UniLog’s performance.
-We will use $UniLog_{{cl}}$ to represent the version based on CodeLlama-7B, $UniLog_{{ds}}$ to represent the version based on DeepSeek-V3.
-FastLog: FastLog (Xie et al., 2023) defines the logging task in two steps: finding the position and generating and inserting a complete log statement into the source code.
+We will use $\bm{UniLog_{{cl}}}$ to represent the version based on CodeLlama-7B, $\bm{UniLog_{{ds}}}$ to represent the version based on DeepSeek-V3.
+FastLog: FastLog (Xie et al., 2024) defines the logging task in two steps: finding the position and generating and inserting a complete log statement into the source code.
 This approach avoids rewriting the source code, a key limitation of $LANCE$.
@@
 Static evaluation on the complete logging task was conducted.
-The values in the Original lines indicate the reported results from the respective methods’ own evaluations.
-The best performance evaluated by AL-Bench across different metrics is highlighted.
-Method PA LA MA DVA STB (BLEU)
-Now Original Now Original Now Original
-FastLog 57.54 $58.84$ 62.72 $59.75$ 6.90 $4.52$ 17.66 20.20
-$UniLog_{{ds}}$ $36.31$ $76.90$ $36.31$ $72.30$ $5.19$ $22.40$ 15.73 11.60
-$UniLog_{{cl}}$ $23.29$ $76.90$ $23.29$ $72.30$ $2.31$ $22.40$ 16.08 8.43
-LANCE $34.67$ $65.40$ $34.67$ $66.24$ $2.99$ $16.90$ 14.42 6.52
-$LEONID_{S}$ $15.22$ $76.45$ $15.22$ $73.53$ $1.38$ $31.55$ 5.35 2.03
-4.2. RQ1: How well are logging tools at predicting log statements in static evaluation?
+Note that lower ALD means better log level quality, and BLEU-4 & ROUGE-L refer to two kinds of STS metrics.
+The best results are highlighted.
+Method PA LA MA ALD DEA BLEU-4 ROUGE-L
+FastLog 58.39 62.15 6.93 0.63 18.00 20.14 29.32
+$UniLog_{{ds}}$ $37.11$ $60.66$ $5.23$ 0.61 16.20 11.62 26.37
+$UniLog_{{cl}}$ $23.49$ $50.97$ $2.71$ 0.79 16.60 8.79 29.88
+LANCE $35.97$ $37.70$ $3.11$ 2.11 15.25 6.7 15.64
+$LEONID_{S}$ $11.26$ $17.90$ $1.96$ 3.78 8.51 3.45 6.69
+4.2. RQ1: How similar are automatically generated log statements to human-written logging practices?
+Table 3.
+The performance of tools when facing different length input data.
+Short means data shorter than 512 tokens, Long means data longer than 512 tokens, and $\Delta$ means the difference in logging tools performance when facing the Short data and the Long data.
+Method Dataset PA LA MA ALD DEA BLEU-4 ROUGE-L
+FastLog Short 62.27 61.79 6.81 0.61 19.65 19.71 29.43
+Long 45.04 63.37 7.37 0.68 12.35 21.63 28.9
+$\Delta$ $\downarrow 17.23$ $\uparrow 1.58$ $\uparrow 1.58$ $\uparrow 0.56$ $\downarrow 7.30$ $\uparrow 1.92$ $\downarrow 0.47$
+$UniLog_{d}{}_{s}$ Short 41.19 59.00 4.11 0.63 16.21 9.93 24.51
+Long 23.10 66.38 9.06 0.54 16.19 17.40 32.77
+$\Delta$ $\downarrow 18.09$ $\uparrow 7.38$ $\uparrow 4.95$ $\downarrow 0.09$ $\downarrow 0.03$ $\uparrow 7.47$ $\uparrow 8.26$
+$UniLog_{c}{}_{l}$ Short 27.78 51.07 2.64 0.77 17.53 8.86 30.24
+Long 8.73 50.59 2.94 0.82 13.39 8.44 24.58
+$\Delta$ $\downarrow 19.05$ $\downarrow 0.48$ $\uparrow 0.30$ $\uparrow 0.09$ $\downarrow 4.14$ $\downarrow 0.42$ $\downarrow 5.66$
+LANCE Short 46.44 48.67 4.01 1.56 19.69 8.65 20.19
+Long 0 0 0 0 0 0 0
+$\Delta$
+$LEONID_{S}$ Short 14.54 23.11 2.53 3.43 10.99 4.45 8.64
+Long 0 0 0 0 0 0 0
+$\Delta$
+We evaluated four automatic logging tools using AL-Bench, leveraging 39,600 static evaluation pairs.
+The evaluation results are presented in Table 2.
+All logging tools show average accuracy drops of 37.49%, 23.43%, and 15.80% in predicting log position, level, and message compared to their reported results.
+Most strikingly, $LEONID_{S}$ collapses from 31.55 → 1.96 in MA ($\downarrow 93.8\%$) and $UniLog_{cl}$ plummets from 76.90 → 23.49 in PA ($\downarrow 69.5\%$).
+Those drops expose severe overfitting to narrow, synthetic datasets in prior studies.
+When evaluated on AL-Bench, state-of-the-art tools achieved limited scores of 18.00 (DEA), 20.14 (BLEU-4), and 29.88 (ROUGE-L) in STS.
+These results highlight persistent limitations in generating semantically coherent natural language descriptions and reliably identifying contextually relevant dynamic content for logging.
+To investigate the performance discrepancy relative to prior studies, we partitioned the dataset into two subsets: instances with sub-512 token lengths and those exceeding 512 tokens.
+This division enables us to determine whether instance length contributes to the performance discrepancies observed among the logging tools, providing further insight into how different data characteristics influence tool effectiveness.
+After dividing the data, we obtained two groups: long data containing 8,925 instances and short data with 30,675 instances.
+The evaluation results in Table 3 show a clear difference in tool performance based on the length of the instances.
+Notably, LANCE and LEONID struggled with instances longer than 512 tokens, failing to generate syntax-correct code and, in some cases, producing incomplete code.
+This explains why their scores for these cases are reported as zero, highlighting the input length limitations of both tools and their inability to handle longer, more complex instances effectively.
+$UniLog_{cl}$ and FastLog show a considerable drop in PA and DEA when handling longer data, indicating that they struggle to predict log positions and select dynamic variables in more complex instances.
+$UniLog_{ds}$ exhibits a significant decline in positional accuracy but only marginal degradation in dynamic variable identification.
+Notably, it demonstrates stronger performance on metrics such as MA and STS when processing longer input sequences.
+This observation implies that enhancing a model’s architectural foundation and providing richer contextual data can improve its capacity to generate syntactically and semantically coherent logging statements through greater contextual awareness.
+Although $UniLog_{ds}$ is equipped with a more powerful backbone, it still struggles to choose the appropriate position for the log statement.
+As widely demonstrated that the LLMs are not good at counting numbers (Ahn et al., 2024), making LLM generate the exact line number might not be a wise choice.
+Furthermore, by simply analyzing the control flow graph of the code, we might be able to exclude positions where logging is not feasible to leave fewer choices.
 Figure 5.
@@
 The performance of each tool varies considerably across different projects and the trends of all methods across all projects generally remain consistent.
-We evaluated four automatic logging tools using the AL-Bench static evaluation dataset and corresponding static evaluation methods, ensuring a fair comparison of their performance.
-Following previous studies (Mastropaolo et al., 2022; Xu et al., 2024; Xie et al., 2023; Mastropaolo et al., 2023), we first processed our dataset to create evaluation pairs $<M_{s},M_{t}>$ with $M_{s}$ representing the input provided to the model (i.e., $M_{s}$ with one removed log statement) and $M_{t}$ being the expected output (i.e., $M_{t}$ is the removed log statement).
-After processing, we got 42,224 instances.
-The evaluation results are presented in Table 2.
-The metrics for Position Accuracy (PA), Level Accuracy (LA), and Message Accuracy (MA) are adapted from the evaluation frameworks used by the selected logging tools.
-Therefore, we label the values of these metrics as reported in their respective evaluation studies.
-The results demonstrate significant performance differences among the four logging tools.
-FastLog consistently outperforms the other tools across all metrics, but it still faces challenges in specific areas.
-For instance, it scores only 17.66 in Dynamic Variable Accuracy (DVA) and 20.44 in Static Text BLEU (STB), suggesting that even the best tool struggles to generate meaningful natural language descriptions and accurately select dynamic variables for logging.
-Despite these shortcomings, FastLog shows only minor deviations from previous studies, maintaining stable performance across all metrics.
-In contrast, the other three tools, particularly LEONID, experience a significant decline in accuracy across all evaluated metrics.
-As illustrated in Figure 7, the performance of each tool varies considerably across different projects.
-This fluctuation suggests that the reliability and stability of logging tools remain inconsistent when applied to diverse scenarios.
-The performance fluctuations of logging tools across different projects are due to variations in project complexity and the diverse logging requirements.
-Additionally, limited training data diversity may hinder the tools’ ability to perform consistently across unfamiliar scenarios.
-These factors collectively contribute to the observed instability in logging tool performance across different projects.
-As demonstrated in our results, the variability in tool performance across different projects underscores the necessity of diverse evaluation data to capture the full spectrum of a tool’s capabilities and limitations.
-To investigate the causes of performance differences compared to previous studies, we observed that earlier evaluation datasets were filtered to exclude instances longer than 512 tokens.
-To assess the impact of this filtering, we divided our dataset into two groups: instances with shorter than 512 tokens and those with longer than 512 tokens.
-This division enables us to determine whether instance length contributes to the performance discrepancies observed among the logging tools, providing further insight into how different data characteristics influence tool effectiveness.
-Table 3.
-The performance of tools when facing different length input data.
-Long means data longer than 512 tokens, Short means data shorter than 512 tokens, and $\Delta$ means the difference in logging tools performance when facing the Short data and the Long data.
-Method Dataset PA LA MA DVA STB
-FastLog Long 43.36 64.92 7.02 11.98 20.97
-Short 61.79 62.33 6.88 19.36 19.98
-$\Delta$ $\downarrow 18.43$ $\uparrow 2.59$ $\uparrow 0.14$ $\downarrow 7.38$ $\uparrow 0.99$
-$UniLog_{d}{}_{s}$ Long 22.28 66.66 8.82 15.56 17.05
-Short 40.51 59.52 4.11 15.78 9.97
-$\Delta$ $\downarrow 18.23$ $\uparrow 7.14$ $\uparrow 4.11$ $\downarrow 0.22$ $\uparrow 7.08$
-$UniLog_{c}{}_{l}$ Long 8.70 51.62 2.84 12.91 8.20
-Short 27.67 51.78 2.62 17.02 8.80
-$\Delta$ $\downarrow 18.97$ $\downarrow 0.16$ $\uparrow 0.22$ $\downarrow 4.12$ $\downarrow 0.60$
-LANCE Long 0 0 0 0 0
-Short 45.05 47.61 3.89 18.74 8.12
-$\Delta$
-$LEONID_{S}$ Long 0 0 0 0 0
-Short 19.78 21.46 1.79 6.95 2.64
-$\Delta$
-After dividing the dataset, we obtained two groups: a longer dataset containing 9,732 instances and a shorter dataset with 32,492 instances.
-The evaluation results in Table 3 show a clear difference in tool performance based on the length of the instances.
-Notably, LANCE and LEONID struggled with instances longer than 512 tokens, failing to generate syntax-correct code and, in some cases, producing incomplete code.
-This explains why their scores for these cases are reported as zero, highlighting the input length limitations of both tools and their inability to handle longer, more complex instances effectively.
-$UniLog_{cl}$ and FastLog show a considerable drop in PA and DVA when handling longer data, indicating that they struggle to predict log positions and select dynamic variables in more complex instances.
-$UniLog_{ds}$also shows a dramatic drop in position accuracy, but only a slight decrease in the ability to decide dynamic variables for recording.
-However, with more sufficient information and a more powerful backbone model, it demonstrates significantly better performance in other metrics.
-This indicates that with a more powerful backbone, UniLog can better understand the context with longer inputs.
-Although $UniLog_{d}s$ is equipped with a more powerful backbone, it still struggles to choose the appropriate position for the log statement.
-As widely demonstrated that the LLMs are not good at counting numbers (Ahn et al., 2024), we might change the output format by directly adding log statements into the source code rather than generating the exact line number.
-Furthermore, by simply analyzing the control flow graph of the code, we might be able to exclude positions where logging is not feasible to leave fewer choices.
+To evaluate tool stability, we analyzed performance consistency across individual projects.
+As shown in Figure 5, tool performance exhibits significant variability across projects.
+For instance, in Flink, the state-of-the-art tool achieves an MA score of 13.68, contrasting sharply with the overall project average of 6.93.
+Similarly, in Keycloak, the state-of-the-art DEA reaches 43.84 compared to the aggregate average of 16.60.
+These stark discrepancies underscore inconsistencies in the reliability and generalizability of logging tools when applied to projects with distinct requirements and contexts.
 Answer to RQ1.
-FastLog performs best in AL-Bench ’s static evaluation but still faces challenges in generating meaningful descriptions and selecting key dynamic variables.
-All tools struggle to maintain stable performance under varying logging requirements and complex data, especially in determining log positions.
-Since LLMs struggle with counting, it might be better to avoid outputting exact line numbers and instead add tags in the source code.
-Additionally, analyzing the control graph to exclude impossible positions might be useful to enhance log position determination.
+All logging tools show average accuracy drops of 37.49%, 23.43%, and 15.80% in predicting log position, level, and message compared to their reported results.
+All logging tools exhibit persistent limitations in generating semantically coherent log descriptions and reliably identifying contextually relevant dynamic variables.
+Moreover, they fail to achieve consistent performance reliability across diverse logging requirements and project-specific contexts.
 Table 4. Compilation Failure Rates Across Methods.
-Methods Compilation Success Rates
+Methods CSR
 FastLog 79.9%
@@
 Table 5. Compilation Failure Reasons Analysis.
-Failed Reason Failed Number
+Failed Reason # Failed
 Using Wrong Logging Name 56
@@
 Total 100
-4.3. RQ2: How well can the predicted log statements be compiled successfully?
-To evaluate the tools’ ability to generate compilable log statements, we replace existing logs with predicted ones in the dynamic evaluation dataset and recompile the project to check for successful compilation.
+4.3. RQ2: How compilable are the log statements generated by automatic logging tools?
+To evaluate the tools’ ability to generate compilable log statements, we replace existing log statements with predicted ones and recompile the project to check for successful compilation.
 As shown in Table 5, the best-performing tool, FastLog, achieves a 79.9% Compilation Success Rate, followed by $UniLog_{c}l$ at 70.3%, $UniLog_{d}s$ at 60.2%, LANCE at 49.4%, $LEONID_{M}$ and $LEONID_{S}$ at 25.0% and 16.4%, respectively.
@@
 In contrast, FastLog and UniLog focus solely on generating the new log statement, minimizing the risk of errors by limiting codebase modifications.
-Although FastLog is one of the best tools available, it still leads to significant instances of compile failure.
-When log statements fail to compile, they can prevent the system from functioning as intended and complicate debugging or error tracking.
-To ensure the reliability of automatic logging tools in real-world environments, it’s crucial that the log statements they generate integrate smoothly into the existing codebase.
-Given the importance of reliable compilation, we conducted a manual review of the compilation failures to identify the specific causes.
-This analysis will provide further insight into the key weaknesses of these tools and potential areas for improvement.
+Although FastLog is one of the best tools available, it still leads to significant instances (20.1%) of compile failure.
+To understand the reason for the compilation failures reason, we conducted a manual review of the compilation failures to identify the specific causes.
 We randomly selected 100 failed instances from the best tool, FastLog, and the first two authors cross-checked the causes.
@@
 Current tools are designed for function-level input, highlighting the need for logging tools to integrate better context awareness and validation checks to ensure compatibility with the existing codebase.
-The less frequent errors, such as incompatible types and unreachable statements, also indicate challenges in generating log statements that integrate well with the surrounding code logic.
-The issue of Unreachable Statements is notable because it points to a fundamental weakness in the current logging tools—specifically, their lack of understanding of the code’s control flow.
+The less frequent errors are also noticeable.
+For instance, the issue of unreachable statements points to a weakness in that tools lack an understanding of the code’s control flow, resulting in log statements placed in non-executable paths.
+Collectively, these errors highlight the need for automated logging tools to adopt enhanced context-aware strategies, improving the accuracy and contextual relevance of generated logs.
 Answer to RQ2.
@@
 To improve the reliability of automatic logging tools, it is crucial that they incorporate mechanisms to gather and utilize this additional context during the log statement generation process.
-4.4. RQ3: How well do the generated log statements print logs in dynamic evaluation?
-To answer RQ3, we used the dynamic evaluation dataset including code snippets paired with corresponding unit tests to generate logs.
-For each instance, we replaced the existing log statements in the source code with the predicted log statements from the automatic logging tools.
-The unit tests were then executed to observe the logs generated by the modified code.
+4.4. RQ3: How consistent are runtime logs from auto-logged programs with oracle-generated logs?
+To answer RQ3, we compare the semantic similarity of logs generated by predicted log statements and by the oracle.
 It is important to highlight that, aside from $LEONID_{M}$, all other tools operate under the strong assumption that the given code snippet requires exactly one log statement.
-To highlight this inappropriate assumption, we only allow the tools one chance to predict the log statement, even in cases where multiple log statements might be needed.
-This experimental setup closely mirrors real-world conditions and allows for a more thorough evaluation of the tool’s ability to handle dynamic and varied logging requirements, moving beyond the oversimplified assumption that each snippet requires only one log statement.
-ultimately limiting tool performance in practical applications.
+To highlight the inappropriateness of this assumption, we only allow the tools one chance to predict the log statement, even in cases where multiple log statements are needed.
+Our experimental design emulates real-world logging scenarios by abandoning the assumption that single log statements universally suffice, thereby exposing a critical limitation in existing tools’ capacity to determine contextually appropriate logging density for a given code context.
 Table 6. The semantic similarity between logs printed by the predicted log statements and the ground truth.
-Method Cosine Similarity BLEU BLEU-1 BLEU-4 ROUGE-1 ROUGE-L
-$FastLog$ 0.213 16.172 24.829 15.963 24.427 23.841
-$UniLog_{{cl}}$ 0.174 13.625 19.015 13.423 19.167 18.749
-$UniLog_{{ds}}$ 0.130 11.607 14.674 11.556 14.002 13.869
-$LANCE$ 0.099 8.201 11.342 7.971 11.235 11.033
-$LEONID_{M}$ 0.072 5.466 7.946 5.303 8.095 7.944
-$LEONID_{S}$ 0.044 3.220 4.908 2.955 5.099 5.012
+Method COS(%) BLEU-1 BLEU-2 BLEU-3 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L
+$FastLog$ 21.32 24.83  19.62  17.49 15.96 24.43  18.74 23.84
+$UniLog_{{cl}}$ 17.38 19.02  15.62  14.34 13.42 19.17  15.37 18.75
+$UniLog_{{ds}}$ 13.04 14.67  12.72  12.02 11.56 14.00  12.23 13.87
+$LANCE$ 9.93 11.34  9.21  8.50 7.97 11.23  8.94 11.03
+$LEONID_{M}$ 7.19 7.95  6.51  5.87 5.30 8.10  6.43 7.94
+$LEONID_{S}$ 4.45 4.91  3.82  2.95 2.96 5.10  3.86 5.01
 Table 6 compares the semantic similarity between logs from source and predicted log statements.
 The results indicate that logs from the predicted statements significantly deviate from the ground truth, with consistently low scores across metrics like Cosine Similarity, BLEU, and ROUGE.
-Specifically, low similarity scores indicate that generated logs frequently fall short of matching expected outputs.
 Limiting each code snippet to a single log statement often sharply compromises log similarity, especially when multiple statements are needed.
@@
 False Positive and False Negative Logging Rates across Methods.
-FPLG means the predicted log statement record logs when it should not, and FNLG means the predicted log statement does not record logs when it should.
-Method FPLG Rate FNLG Rate
+FPLR means the predicted log statement record logs when it should not, and FNLR means the predicted log statement does not record logs when it should.
+Method FPLR FNLR
 FastLog 9.28% 18.28%
@@
 $LEONID_{S}$ 8.15% 8.69%
-$LEONID_{M}$ 7.32% 11.6%
-Table 8. FPLG and FNLG Reason Analysis. For FNLG, the major reason is beyond the execution path, and for FPLG, the major reason is lower verbosity level.
-Situation Reason Number
-FNLG Beyond Execution Path 35
+$LEONID_{M}$ 7.32% 11.60%
+Table 8. FPLR and FNLR Reason Analysis. For FNLR, the major reason is beyond the execution path, and for FPLR, the major reason is lower verbosity level.
+Situation Reason # False
+FNLR Beyond Execution Path 35
 Lower Verbosity Level 24
 Wrong Code Format 4
-FPLG Higher Verbosity Level 30
+FPLR Higher Verbosity Level 30
 Beyond Execution Path 3
@@
 To understand the low semantic similarity scores, we examined the log generation process and found many instances where predicted logs were either redundant or missed key information present in the original logs.
-We quantify this issue using two metrics: False Positive Log Generation (FPLG) and False Negative Log Generation (FNLG), as reported in Table 8.
-For example, FastLog reports a 9.2% False Positive Log Rate (FPLR), meaning logs record redundant information in 9.2% of cases, and an 18.28% False Negative Log Rate (FNLR), indicating expected information in logs were missing in 18.28% of cases.
+We quantify this issue using two metrics: FPLR and FNLR, as reported in Table 8.
+For example, FastLog reports a 9.28% FPLR, meaning logs record redundant information in 9.28% of cases, and an 18.28% FNLR, indicating expected information in logs was missing in 18.28% of cases.
 We sampled 100 examples from FastLog, the state-of-the-art (SOTA) model, and manually analyzed the reasons for mismatches with the original logs.
 The results are presented in Table 8.
-We found that the primary reasons for failures in FNLG and FPLG differ significantly.
-For FNLG, the most common issue was the predicted log statements being beyond the execution path (35 cases), followed by lower verbosity levels (24 cases), and a smaller number caused by wrong code format (4 cases).
-In contrast, for FPLG, the main problem was higher verbosity levels (30 cases), with a few instances of log statements being beyond the execution path (3 cases).
+We found that the primary reasons for failures in FNLR and FPLR differ significantly.
+For FNLR, the most common issue was the predicted log statements being beyond the execution path (35 cases), followed by lower verbosity levels (24 cases), and a smaller number caused by wrong code format (4 cases).
+In contrast, for FPLR, the main problem was higher verbosity levels (30 cases), with a few instances of log statements being beyond the execution path (3 cases).
 Overall, verbosity mismatches and execution path discrepancies were the dominant contributors, highlighting challenges in aligning predicted logs with actual logging requirements.
-The factors leading to FPLG and FNLG underscore a critical issue: while static metrics offer valuable insights into the quality of generated log statements, the actual logs are shaped by numerous contextual factors.
+The factors leading to FPLR and FNLR underscore a critical issue: while static metrics offer valuable insights into the quality of generated log statements, the actual logs are shaped by numerous contextual factors.
 Without a thorough understanding of the execution context, it is not possible to comprehensively evaluate the quality of log statements.
@@
 This issue majorly arises from the tools’ lack of awareness of verbosity thresholds and the control graph of the code, which limits their ability to adjust verbosity and determine appropriate log positions based on the context or execution requirements.
-These two limitations highlight that current logging tools lack the adaptability and context-awareness needed for effective real-world application.
+These two limitations highlight that current logging tools lack the adaptability and context-awareness needed for effective real-world applications.
 Answer to RQ3.
-The best predicted log statements by FastLog achieve only 0.231 cosine similarity with the original logs.
+The best-predicted log statements by FastLog achieve only 21.32% cosine similarity with the original logs.
 Many predictions record redundant information, while others miss key details.
@@
 5. Threats to Vadility
-  • Construct Validity: We use BLEU to assess the quality of the generated log messages.
-Although text similarity metrics may not fully capture the quality of the generated text (Gros et al., 2020; Roy et al., 2021), we follow previous works (Xie et al., 2023; Xu et al., 2024; He et al., 2018a; Ding et al., 2023) to use BLEU as widely-accepted quality measures for the generated log messages.
-We also adapt Cosine Similarity based on TF-IDF to evaluate the quality of log content.
-Since log content, excluding the log header, is primarily composed of natural language, cosine similarity effectively captures the semantic meaning (Salton et al., 1975).
+  • Construct Validity: The first threat comes from the use of two widely-accepted syntactic metrics in previous logging research (BLEU and ROUGE) (Xie et al., 2024; Xu et al., 2024a; He et al., 2018a; Ding et al., 2023), as they may not fully capture the semantics of the log messages (Gros et al., 2020; Roy et al., 2021).
+To mitigate this threat, we also adapt Cosine Similarity to evaluate the semantic quality of log messages.
+Since all log headers are excluded, leaving log messages primarily in natural language, Cosine Similarity can effectively capture the log’s semantic meaning (Salton et al., 1975).
   • Internal Validity: The second threat to validity concerns reproducing the baseline.
-To minimize inconsistencies, we adapted the released models from previous work (Xie et al., 2023; Mastropaolo et al., 2022, 2023) and sought guidance from the authors of the closed-source model UniLog (Xu et al., 2024) to reproduce the tool under supervision.
-  • External Validity: In this work, we evaluate the tools with 10 widely recognized Java projects in different domains.
-The results of each tool are examined in different requirements of logging practices.
-The novel evaluation method we propose is not limited to Java projects, we will explore its effects in other programming languages in the future.
+To minimize inconsistencies, we adapted the released models from previous work (Xie et al., 2024; Mastropaolo et al., 2022, 2023) and sought guidance from the authors of the closed-source model UniLog (Xu et al., 2024a) to reproduce the tool under supervision.
+  • External Validity: The third threat comes from the potential data contamination risks in our dataset.
+To mitigate it, we follow the methodology used by GPT-3 (Brown et al., 2020) to assess data contamination severity, i.e., considering test samples with more than a 13-gram overlap with the original code samples as contaminated.
+Specifically, we perform a 13-gram overlap analysis between test and training samples of all evaluation methods.
+Results indicate a peak contamination rate of 1.24% across all training datasets, empirically demonstrating no evidence of data leakage in our framework.
 6. Conclusion
-This paper introduces AL-Bench, designed specifically for automatic logging tools.
-It includes a high-quality dataset and a novel dynamic evaluation method focused on runtime logs, addressing key limitations of prior studies and bridging the gap between real-world requirements and existing evaluation frameworks.
-This dynamic evaluation method assesses both the compilability of predicted log statements and their effectiveness in generating runtime logs.
-Using AL-Bench, we evaluate popular end-to-end automatic logging tools and find that generated log statements fail to compile in 20.1% to 83.6% of cases.
-Even the best predictions achieve only 0.213 cosine similarity between generated and ground-truth runtime logs.
-These results show that automatic logging still has a long way to go.
+In this paper, we introduced AL-Bench, a comprehensive benchmark for automatic logging tools evaluation, featuring a diverse dataset from 10 established projects and a novel dynamic evaluation methodology.
+Our benchmark revealed significant limitations in existing tools through both static and dynamic perspectives: substantial accuracy drops in log position, level, and message prediction compared to reported results, high compilation failure rates (20.1%-83.6%), and low runtime log similarity (maximum 21.32%).
+These findings highlight critical gaps between current tool capabilities and practical requirements.
+We believe AL-Bench establishes a solid foundation for advancing automatic logging research by providing standardized evaluation metrics and exposing areas requiring improvement.
 7. DATA AVAILABILITY